### v1

# Permutations, substitutions and finite axiomatizability

Abstract.
Algebras of relations form an algebraic framework for the study of logical systems, extending the correspondence between Boolean algebras and propositional logic. Tarski’s representable cylindric algebras $\mathsf{RCA}_{\alpha}$, and Halmos’ representable polyadic algebras $\mathsf{RPA}_{\alpha}$ both provide algebraic counterparts to first-order logic. In this paper, we show that the usual finite set of polyadic axioms axiomatize $\mathsf{RPA}_{\alpha}$ over $\mathsf{Rdf}_{\alpha}$, the diagonal-free subreducts of elements in $\mathsf{RCA}_{\alpha}$. In short: $\mathsf{RPA}_{\alpha}=\mathsf{PA}_{\alpha}+\mathsf{Rdf}_{\alpha}$.
Key words and phrases:
Algebraic logic, representable polyadic algebra, substitution-cylindric algebra, cylindric algebra, canonical variety, finite axiomatization, finite-variable logic
1991 Mathematics Subject Classification:
Primary 03G15; Secondary 03E25, 03C05, 03B10
1. What was and what is to come
Algebras of relations have played an important role in the development of logic and its surrounding fields ever since the pioneering works of De Morgan, Peirce, and Schröder initiated the area. Boolean set algebras (that is, algebras of unary relations) are the algebraic counterparts of propositional logic: the equations valid in Boolean set algebras correspond to propositional tautologies. Alfred Tarski and Paul Halmos introduced the classes $\mathsf{RCA}_{\alpha},\mathsf{Rdf}_{\alpha}$ and $\mathsf{RPEA}_{\alpha},\mathsf{RPA}_{\alpha}$ of algebras isomorphic to cylindric and polyadic algebras of $\alpha$-ary relations, respectively, as candidates for playing the same role in first-order logic (with and without equality) using $\alpha$ many variables as Boolean set algebras play in propositional logic. The class $\mathsf{BA}$ of Boolean algebras, the class of all algebras isomorphic to Boolean set algebras, is axiomatized by finitely many equations. Similarly, all of the above-mentioned classes of cylindric and polyadic algebras are axiomatizable by equations if $\alpha$ is finite. However, unlike $\mathsf{BA}$, none of them can be axiomatized by finitely many equations if $\alpha\geq 3$. [^1] For $\mathsf{RCA}_{\alpha}$, non-finite axiomatizability is proved by J. Donald Monk [16], and for the rest by James Johnson [13]. In contrast, for $\alpha<3$, all of them have finite axiom systems, see [12, 5.1.77, 3.2.65, 5.4.33, p.242].
From now on, throughout the paper, $\alpha\geq 3$ is a finite ordinal. $T_{\alpha}$, or just $T$, denotes the set of all transformations $\sigma:\alpha\to\alpha$.
Cylindric and polyadic algebras are Boolean algebras with some extra operations. The similarity type of $\mathsf{RPEA}_{\alpha}$ has the following extra Boolean operations: the unary cylindrification operations $\mathsf{c}_{(\Gamma)}$ and substitution operations $\mathsf{s}_{\sigma}$ for all subsets $\Gamma\subseteq\alpha$ and transformations $\sigma\in T$ together with the diagonal constants $\mathsf{d}_{ij}$ for all $i,j<\alpha$. The other classes are obtained by discarding some of these operations and taking subalgebras: $\mathsf{RPA}_{\alpha}$ is the class of subalgebras of reducts of elements of $\mathsf{RPEA}_{\alpha}$ when discarding the diagonal constants $\mathsf{d}_{ij},i,j<\alpha$; $\mathsf{RCA}_{\alpha}$ is the class of subalgebras of reducts of elements of $\mathsf{RPEA}_{\alpha}$ when discarding the substitution operations $\mathsf{s}_{\sigma},\ \sigma\in T$; and $\mathsf{Rdf}_{\alpha}$ is the class of subalgebras of elements of $\mathsf{RPEA}_{\alpha}$ when we keep only the cylindrification operations. Hence, $\mathsf{Rdf}_{\alpha}$ is the class of all subalgebras of elements of $\mathsf{RPA}_{\alpha}$ with the substitution operations discarded. See Figure 1.
Which of the extra Boolean operations introduce infinitely many new axioms, thereby causing non-finite axiomatizability? The cylindrifications clearly bring in infinitely many new axioms, since $\mathsf{Rdf}_{\alpha}$ is non-finitely axiomatizable by [13] while $\mathsf{BA}$ is finitely axiomatizable. A natural question arises whether only the cylindrification operations cause non-finite axiomatizability, or do the diagonal constants and the substitution operations contribute, too. Whether the diagonal constants or the substitution operations bring in infinitely many new axioms is asked in [13]. Namely, Problems 1,2 in [13] ask whether $\mathsf{RCA}_{\alpha}$ is finitely axiomatizable over $\mathsf{Rdf}_{\alpha}$, and whether $\mathsf{RPEA}_{\alpha}$ is finitely axiomatizable over $\mathsf{RPA}_{\alpha}$ or over $\mathsf{RCA}_{\alpha}$. [^2] If $K$ is a class of structures of similar types and $K^{\prime}$ is a class of structures whose type is a reduct of the type of $K$, we say that $K$ is finitely axiomatizable over $K^{\prime}$ if there is a finite set of axioms $\Gamma$ such that $\mathfrak{A}\in K$ iff $\mathfrak{A}$ is a model of $\Gamma$ and the reduct of $\mathfrak{A}$ to the type of $K^{\prime}$ is a member of $K^{\prime}$. All these questions have been answered negatively in the meantime [^3] In [1, Thm.4], [1, Thm.6], and [5], respectively., providing the answer to the question that all three kinds of extra-Boolean operations contribute to non-finite axiomatizability.
$\mathsf{RPEA}_{\alpha}$$\mathsf{RCA}_{\alpha}$$\mathsf{RPA}_{\alpha}$$\mathsf{Rdf}_{\alpha}$$\mathsf{s}_{\sigma}$$\mathsf{d}_{ij}$$\mathsf{d}_{ij}$$\mathsf{s}_{\sigma}$
Figure 1. The three thick arrows represent relative non-finite axiomatizability. In the fourth line there is relative finite axiomatizability, this is proved in the present paper.
One natural question, however, was not asked in [13]. This is whether $\mathsf{RPA}_{\alpha}$ is finitely axiomatizable over $\mathsf{Rdf}_{\alpha}$ or not. This question is answered in the present paper. Surprisingly, we get an affirmative answer: $\mathsf{RPA}_{\alpha}$ is finitely axiomatizable over $\mathsf{Rdf}_{\alpha}$. This is in sharp contrast with the situation when the diagonal constants are available: $\mathsf{RPEA}_{\alpha}$ can be axiomatized over $\mathsf{RCA}_{\alpha}$ with infinite and rather complex system of equations only (see the main theorem of [5]). Thus, we have to refine the picture: the substitution operations $\mathsf{s}_{\sigma}$ contribute to non-finite axiomatizability only when the diagonal constants are present in the similarity type, otherwise they do not contribute to non-finite axiomatizability.
Since all of $\mathsf{RPEA}_{\alpha},\mathsf{RPA}_{\alpha},\mathsf{RCA}_{\alpha},\mathsf{Rdf}_{\alpha}$ are nonfinitely axiomatizable, finitely axiomatized “approximating” classes $\mathsf{PEA}_{\alpha},\mathsf{PA}_{\alpha},\mathsf{CA}_{\alpha},\mathsf{Df}_{\alpha}$ were introduced for their study. In these varieties, the most important properties of the operations were gathered as axioms, and in such situations the question whether the finite axioms were well selected always is in the air. We prove in this paper that the finitely many polyadic axioms introduced by Paul Halmos in the definition of $\mathsf{PA}_{\alpha}$ to describe the behaviour of the substitution operations [^4] See, for example, [10, pp.27–28], and [12, Def.5.4.1]. are adequate in the sense that the equations (1)-(4) below together with the equational theory of $\mathsf{Rdf}_{\alpha}$ axiomatize the equational theory of $\mathsf{RPA}_{\alpha}$:
Theorem 1.1 ($\mathsf{s}_{\sigma}$ is finitely axiomatizable over $\mathsf{c}_{(\Gamma)}$).
Assume that $3\leq\alpha$ is finite. Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{(\Gamma)},\mathsf{s}_{\sigma}\rangle_{\Gamma\subseteq\alpha,\sigma\in T}$. Then $\mathfrak{A}\in\mathsf{RPA}_{\alpha}$ if and only if $\langle A,+,-,\mathsf{c}_{(\Gamma)}\rangle_{\Gamma\subseteq\alpha}\in\mathsf{Rdf}_{\alpha}$ and the following equations hold in $\mathfrak{A}$ for all $\Gamma\subseteq\alpha$ and $\sigma,\tau\in T$.
$\mathsf{s}_{\sigma}(x+y)=\mathsf{s}_{\sigma}(x)+\mathsf{s}_{\sigma}(y),\quad\mathsf{s}_{\sigma}(-x)=-\mathsf{s}_{\sigma}(x),$ (1)
$\mathsf{s}_{\sigma\circ\tau}(x)=\mathsf{s}_{\sigma}\mathsf{s}_{\tau}(x),\quad\mathsf{s}_{Id}(x)=x,$ (2)
$\mathsf{s}_{\sigma}\mathsf{c}_{(\Gamma)}x=\mathsf{s}_{\tau}\mathsf{c}_{(\Gamma)}x\quad\mbox{if $\sigma,\tau$ differ only in $\Gamma$},$ (3)
$\mathsf{c}_{(\Gamma)}\mathsf{s}_{\sigma}(x)=\mathsf{s}_{\sigma}\mathsf{c}_{(\sigma^{-1}\Gamma)}x,\quad\mbox{if $\sigma$ is one-one on $\sigma^{-1}\Gamma$}.$ (4)
This is one of the main theorems of the present paper. We get Theorem 1.1 as a corollary to a more refined investigation, see Theorems 1.2 and 1.3 below. Let $i/\penalty 50j]\in T$ and $i,j]\in T$ denote the replacement and transposition operations, respectively, where $i,j<\alpha$. They are defined as, for every $k<\alpha$,
[…]
It is known that $T$ is generated by the replacements and transpositions, in particular, each non-surjective transformation is the composition of some replacements, and each surjective transformation (i.e., permutation) is a composition of some transpositions. (Recall that $\alpha$ is finite.)
[…]
Theorem 1.2 ($\mathsf{s}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i}$).
Assume that $3\leq\alpha$ is finite. Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}\rangle_{i,j<\alpha}$. Then $\mathfrak{A}\in\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}$ if and only if $\langle A,+,-,\mathsf{c}_{i}\rangle_{i<\alpha}\in\mathsf{Rdf}_{\alpha}$ and the following equations hold in $\mathfrak{A}$ for all distinct $i,j,k<\alpha$.
$\mathsf{s}_{ij}(x+y)=\mathsf{s}_{ij}(x)+\mathsf{s}_{ij}(y),\quad\mathsf{s}_{ij}(-x)=-\mathsf{s}_{ij}(x),$ (6)
$\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x,\quad\mathsf{s}_{ii}(x)=x,$ (7)
$\mathsf{s}_{ij}\mathsf{c}_{k}x=\mathsf{c}_{k}\mathsf{s}_{ij}x,$ (8)
$\mathsf{s}_{ij}x=\mathsf{c}_{i}\mathsf{s}_{ij}x.$ (9)
Theorem 1.3 ($\mathsf{p}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i},\mathsf{s}_{ij}$).
Assume that $3\leq\alpha$ is finite. Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij},\mathsf{p}_{ij}\rangle_{i,j<\alpha}$. Then $\mathfrak{A}\in\mathsf{RFPA}_{\alpha}$ if and only if $\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}\rangle_{i,j<\alpha}\in\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}$ and the following equations hold in $\mathfrak{A}$ for all distinct $i,j,k<\alpha$.
$\mathsf{p}_{ij}(x+y)=\mathsf{p}_{ij}(x)+\mathsf{p}_{ij}(y),\quad\mathsf{p}_{ij}(-x)=-\mathsf{p}_{ij}(x),$ (10)
$\mathsf{p}_{ij}\mathsf{p}_{ij}x=x,\quad\mathsf{p}_{ii}x=x,\quad\mathsf{p}_{ij}x=\mathsf{p}_{ji}x,$ (11)
$\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x,$ (12)
$\mathsf{p}_{ij}\mathsf{s}_{ij}x=\mathsf{s}_{ji}\mathsf{p}_{ij}x.$ (13)
We do not know whether $\mathsf{p}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i}$, but we know that $\mathsf{s}_{ij}$ is finitely axiomatisable over $\mathsf{c}_{i},\mathsf{p}_{ij}$, see [3].
The question of which of the extra Bolean operations of $\mathsf{RFPEA}_{\alpha}$ bring in infinitely many new equations is systematically investigated in [1]. The answer to this question depends on the order in which the operations are added to our algebras; in [1, Figure 1] the known answers as of 1991 are indicated on the edges of a tree. Theorems 1.2 and 1.3 above provide answers for two edges on that tree marked as unknown. As of now, only one answer remains unknown, namely whether $\mathsf{p}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i}$. [^5] Five edges are marked in [1, Figure 1] as unknown. Two of them are answered in the present paper, one was answerd by Monk [18] in 1997, and another one in the recent [3].
With these two lines filled out, an interesting behaviour of the substitution operations $\mathsf{s}_{ij}$ can be read off the tree: these operations never bring in infinitely many new equations. In contrast, the permutation operations $\mathsf{p}_{ij}$ exhibit a mixed behaviour: they bring in infinitely many new axioms in the presence of the diagonal constants and the cylindrifications, and when $\mathsf{s}_{ij}$ are available but the diagonals are not, they do not contribute to infinite axiomatizability.
The substitution operations $\mathsf{s}_{ij}$ behave rather interestingly in other respects, too, and they are quite well investigated. [^6] See, for example [6], [12, sections 1.5, 1.11], [19], [20], [21], [23], [24]. For example, Charles Pinter proved in [19] that in certain complete algebras cylindrifications $\mathsf{c}_{i}$ and diagonals $\mathsf{d}_{ij}$ can be defined from the $\mathsf{s}_{ij}$ so that these defined operations satisfy the defining equations of $\mathsf{CA}_{\alpha}$. This fact will be one of the key steps of our proofs. Because these intimate connections between $\mathsf{s}_{ij}$ and the $\mathsf{c}_{i},\mathsf{d}_{ij}$ hold only in complete algebras, we will rely on the existence of the canonical embedding algebras, because they are complete. Robert Goldblatt’s results on canonical varieties come to our aid here. We note that Goldblatt has several papers where he applies the duality between structures and complex algebras to algebras of relations, see [7, 8, 9, 2].
Originally, Halmos defined polyadic algebras for algebraizing first-order logic without equality, while Tarski defined cylindric algebras for algebraizing first-order logic with equality. The equality atomic formulas in first-order logic are algebraized by the diagonal constants $\mathsf{d}_{ij}$. Substituting variables in formulas is treated in polyadic algebras explicitly via the substitution operations $\mathsf{s}_{\sigma}$, while they are treated implicitly via defined terms in cylindric algebras, see [12, sec.1.5, end of sec.1.11]. Halmos [10, p.28] raises the problem of understanding the relation between these two treatments of equality and variable-substitutions. The results in the present paper contribute to this investigation. Detailed connections between variable-substitution in formulas and the algebraic substitution operations can be found in [5] and in [3]. It turns out that already on the syntactic level of logic, there is a striking difference between transposition of variables ($\mathsf{p}_{ij}$) and replacement of variables ($\mathsf{s}_{ij}$). For more on consequences of these axiomatizability investigations in first-order logic see Monk [17], Johnson [14] and [5, Sections 1,9].
The paper is organized as follows. In Section 2, we recall the definitions of the main classes of algebras relevant to our study. Section 3 shows that our classes are canonical, that is, they are closed under taking canonical embedding algebras. In Section 4 we introduce equational axiom systems. Then, in Section 5, we prove Theorem 1.2, and in Section 6 we prove Theorems 1.3, 1.1.
2. Setting the stage
In this section, we introduce $\mathsf{Rdf}_{\alpha}$ and $\mathsf{RPA}_{\alpha}$, together with some algebras of $\alpha$-ary relations that we will deal with in the present paper.
We will use that an ordinal $\alpha$ is the set of smaller ordinals, so $\alpha=\{i:i<\alpha\}$, and so a $U$-termed $\alpha$-sequence $s=\langle s_{i}\rangle_{i<\alpha}$ is a function $s:\alpha\to U$ defined by $s(i)=s_{i}$ for all $i<\alpha$.
The set of all $U$-termed $\alpha$-sequences is denoted by $^{\alpha}U$. For a sequence $s\in{}^{\alpha}U$, $i<\alpha$, and $u\in U$ we denote by $s(i/u)$ the sequence we obtain from $s$ by replacing its $i$th value with $u$. We note that $s(i/s(j))=s\circ[i/j]$, and $s(i/s(j))(j/s(i))=s\circ[i,j]$ where $\circ$ denotes function composition. An $\alpha$-ary relation over $U$ is a subset $R\subseteq{}^{\alpha}U$. The following unary operations on $\alpha$-ary relations over $U$ are defined:
$\operatorname{C}_{i}^{U}(R)$ $=\{s\in{}^{\alpha}U:\;s(i/u)\in R\text{ for some }u\in U\},$ (14)
$\mathop{S}\nolimits_{ij}^{U}(R)$ $=\{s\in{}^{\alpha}U:\;s(i/\penalty 50s(j))\in R\},$ (15)
$\mathop{P}\nolimits_{ij}^{U}(R)$ $=\big\{s\in{}^{\alpha}U:\;s(i/s(j))(j/s(i))\in R\big\},$ (16)
$\operatorname{D}_{ij}^{U}$ $=\{s\in{}^{\alpha}U:\;s(i)=s(j)\}.$ (17)
We might omit the superscript $U$ when this is not likely to lead to confusion. We note that $\mathop{P}\nolimits_{ij}^{U}(R)$ does not, in fact, depend on $U$ (nevertheless, for clarity, we often indicate it), while $\operatorname{C}_{i}^{U}(R)$, $\mathop{S}\nolimits_{ij}^{U}(R)$ and $\operatorname{D}_{ij}^{U}$ do depend on $U$. The set of all subsets of a set $V$ is denoted by $\operatorname{Sb}(V)$, so $\operatorname{Sb}({}^{\alpha}U)$ is the set of all $\alpha$-ary relations on $U$.
The $\alpha$-dimensional full finitary polyadic set algebra is the structure
$\mathfrak{A}=\langle\operatorname{Sb}({}^{\alpha}U),\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U},\mathop{P}\nolimits_{ij}^{U}\rangle_{i,j<\alpha}\,.$ (18)
For a class $K$ of algebras, $\mathbf{I}K,\mathbf{S}K,\mathbf{P}K,\mathbf{Up}K$ denote the classes of all isomorphic copies, all subalgebras, and all isomorphic copies of products and of ultraproducts of elements of $K$, respectively. We will focus on the following classes of algebras
$\mathsf{RA}_{\alpha}^{csp}$ $=\mathbf{SI}\big\{\langle\operatorname{Sb}({}^{\alpha}U),\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U},\mathop{P}\nolimits_{ij}^{U}\rangle_{i,j<\alpha}:\;U\text{ is a set}\big\},$ (19)
$\mathsf{RA}_{\alpha}^{cs}$ $=\mathbf{SI}\big\{\langle\operatorname{Sb}({}^{\alpha}U),\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U}\rangle_{i,j<\alpha}:\;U\text{ is a set}\big\},$ (20)
$\mathsf{RA}_{\alpha}^{c}$ $=\mathbf{SI}\big\{\langle\operatorname{Sb}({}^{\alpha}U),\cup,\smallsetminus,\operatorname{C}_{i}^{U}\rangle_{i,j<\alpha}:\;U\text{ is a set}\big\}.$ (21)
The superscripts csp, cs and c refer to the similarity types of the algebras:
$c$ $=\text{Boolean}+\{\mathsf{c}_{i}:\;i,j<\alpha\},$
$cs$ $=\text{Boolean}+\{\mathsf{c}_{i},\mathsf{s}_{ij}:\;i,j<\alpha\},$
$csp$ $=\text{Boolean}+\{\mathsf{c}_{i},\mathsf{s}_{ij},\mathsf{p}_{ij}:\;i,j<\alpha\}.$
If $\mathfrak{A}$ is a $cs$-type algebra, then $\operatorname{Rd}_{c}\mathfrak{A}$ denotes the $c$-type reduct of $\mathfrak{A}$, etc.
Let $\tau$ denote one of c, cs, csp. It is known that $\mathsf{RA}_{\alpha}^{\tau}$ is a discriminator class and it is closed under taking ultraproducts, see e.g., [4, Lemmas 7.1.6 and 7.1.7]. Hence, $\mathsf{RA}_{\alpha}^{\tau}=\mathbf{SUp}\mathsf{RA}_{\alpha}^{\tau}$ is a class axiomatizable by a set of quantifier-free formulas and $\mathbf{SP}\mathsf{RA}_{\alpha}^{\tau}=\mathbf{HSP}\mathsf{RA}_{\alpha}^{\tau}$ is a class axiomatizable by equations, in other words, $\mathbf{SP}\mathsf{RA}^{\tau}_{\alpha}$ is the variety generated by $\mathsf{RA}_{\alpha}^{\tau}$. Subalgebras of reducts of full finitary polyadic set algebras are called set algebras, so $\mathsf{RA}_{\alpha}^{\tau}$ is the class of the $\tau$-type set algebras, up to isomorphisms [^7] In the following, we sometime omit “up to isomorphisms” when this is not likely to cause confusion.. The elements of $\mathbf{SP}\mathsf{RA}_{\alpha}^{\tau}$ are called representable algebras. These are natural algebras of $\alpha$-ary relations. In the literature, $\mathsf{RA}^{c}_{\alpha}$ is called the class of $\alpha$-dimensional diagonal-free cylindric set algebras [12, 5.1], and $\mathsf{RA}^{csp}_{\alpha}$ is the class of $\alpha$-dimensional finitary polyadic set algebras. The classes $\mathsf{Rdf}_{\alpha}$ and $\mathsf{RFPA}_{\alpha}$ of representable diagonal-free and representable finitary polyadic algebras are $\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and $\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$, respectively. Recall that finitary polyadic algebras are term-definitionally equivalent to polyadic algebras as defined by Halmos; let $\mathsf{RPA}_{\alpha}$ denote the class of the latter algebras.
The goal of this paper is to prove that $\mathsf{RPA}_{\alpha}$ is finitely axiomatizable over $\mathsf{Rdf}_{\alpha}$ for finite $\alpha\geq 3$. We will show this by showing that $\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$, and $\mathsf{RA}_{\alpha}^{csp}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{cs}$, and then using term-definability as proved in [22].
3. Canonicity of $\mathsf{RA}^{c}_{\alpha}$ and $\mathsf{RA}^{cs}_{\alpha}$
In this section, we use the notion of canonical embedding algebras, cf. [12, 2.7], [11, 2.7.2, 2.7.3], [4, Thms. 2.8.9, 2.8.15]. First, we recall the basic definitions customized for $c$-type algebras.
Definition 3.1.
A ($c$-type) atom structure of dimension $\alpha$ is a first-order relational structure $\mathcal{S}=\langle X,T_{i}\rangle_{i<\alpha}$, where $T_{i}\subseteq{}^{2}X$ for every $i<\alpha$.
Definition 3.2.
The full complex algebra of an atom structure $\mathcal{S}=\langle X$, $T_{i}$ $\rangle_{i<\alpha}$ is the algebra
$\mathfrak{Cm}(\mathcal{S})=\langle\operatorname{Sb}(X),T_{i}^{*}\rangle,$ (22)
where for any $A\subseteq X$ we let
$T_{i}^{*}(A)=\{y\in X:\;\langle x,y\rangle\in T_{i}\text{ for some }x\in A\}\,.$ (23)
Definition 3.3.
Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i}\rangle_{i<\alpha}$ be a $c$-type algebra, such that $\langle A,+,$-$\rangle$ is a Boolean algebra. The ultrafilter frame of $\mathfrak{A}$ is the first-order relational structure
$\mathfrak{Uf}(\mathfrak{A})=\langle{U\!f}(\mathfrak{A}),T_{i}\rangle_{i<\alpha},$ (24)
where ${U\!f}(\mathfrak{A})$ is the set of ultrafilters of $\mathfrak{A}$, and $T_{i}\subseteq{}^{2}{U\!f}(\mathfrak{A})$ is defined as
$\langle p,q\rangle\in T_{i}$ $\quad\Longleftrightarrow\quad\{\mathsf{c}_{i}(x):\;x\in p\}\subseteq q.$ (25)
Definition 3.4.
$\mathfrak{A}$ is called a Boolean algebra with operators ($\mathsf{BAO}$) if the Boolean reduct of $\mathfrak{A}$ is a Boolean algebra, and the extra Boolean operations are additive. $\mathfrak{A}$ is called a complete BAO if $\mathfrak{A}$ is a BAO, arbitrary suprema exist in the Boolean reduct of $\mathfrak{A}$ and the extra Boolean operations preserve infinite sums argumentwise. $\mathfrak{A}$ is called a union-complete BAO if $\mathfrak{A}$ is a complete BAO with Boolean reduct a set Boolean algebra closed under arbitrary unions. This implies that $\sum X=\bigcup X$ for all nonempty $X$.
The canonical embedding algebra of a $\mathsf{BAO}$ $\mathfrak{A}$, denoted by $\mathfrak{Em}(\mathfrak{A})$, is the full complex algebra of its ultrafilter frame
$\mathfrak{Em}(\mathfrak{A})=\mathfrak{Cm}\mathfrak{Uf}(\mathfrak{A}).$ (26)
By Jónsson–Tarski [15], every $\mathsf{BAO}$ $\mathfrak{A}$ embeds into $\mathfrak{Em}(\mathfrak{A})$, and $\mathfrak{Em}(\mathfrak{A})$ is a union-complete BAO. Another important property of the canonical embedding algebras $\mathfrak{Em}(\mathfrak{A})$ is that they preserve positive equations. A positive equation is an equation in the language of $\mathfrak{A}$ in which negation - does not occur, but the derived operations $\cdot,0,1$ can occur. Thus, $\mathfrak{Em}(\mathfrak{A})\vDash e$ if $\mathfrak{A}\vDash e$, for any positive equation $e$. See, e.g., [4, 2.8.9, 2.8.15], [12, 2.7.5, 2l7.14].
A class $K$ of algebras is called canonical if $\mathfrak{Em}(\mathfrak{A})\in K$ whenever $\mathfrak{A}\in K$.
The following theorem is a corollary of Robert Goldblatt’s results on the duality between structures and Boolean algebras with operators.
Theorem 3.5.
$\mathsf{RA}_{\alpha}^{\tau}$ as well as $\mathbf{SP}\mathsf{RA}_{\alpha}^{\tau}$ are canonical, for $\tau\in\{c,cs,csp\}$.
Proof . We elaborate the proof for $\tau=c$, the proofs for the other cases can be obtained by making obvious modifications. A c-type sequence-structure is
$\langle{}^{\alpha}U,T_{i}\rangle_{i<\alpha}$ (27)
where $U$ is a non-empty set and $sT_{i}z$ iff $z=s(i/\penalty 50u)$ for some $u\in U$. Let $\mathsf{Str}^{c}$ denote the class of all $c$-type sequence-structures. By definition,
$\mathsf{RA}^{c}_{\alpha}=\mathbf{SI}\{\mathfrak{Cm}(\mathcal{S}):\mathcal{S}\in\mathsf{Str}^{c}\}.$ (28)
Now, $\mathsf{Str}^{c}$ is closed under taking ultraproducts by [9, Lemma 3.5]. Hence, $\mathsf{RA}^{c}_{\alpha}$ and $\mathbf{SP}\mathsf{RA}^{c}_{\alpha}$ are canonical, by Corollary 3.6.3 and Theorem 3.6.7 of [7]. $\blacksquare$
All the above, e.g., the definitions of atom structure, full complex algebra, and ultrafilter frame can be extended in a natural way to $cs$-type and cspcsp-type algebras. For completeness, we briefly include the definitions for $cs$-type algebras.
Definition 3.6.
A ($cs$-type) atom structure of dimension $\alpha$ is a first-order relational structure $\mathcal{S}=\langle X,T_{i},R_{ij}\rangle_{i,j<\alpha}$, where $T_{i},R_{ij}\subseteq{}^{2}X$ for every $i,j<\alpha$.
Definition 3.7.
The full complex algebra of a $cs$-type atom structure $\mathcal{S}=\langle X$, $T_{i}$, $R_{ij}$ $\rangle_{i,j<\alpha}$ is the algebra
$\mathfrak{Cm}(\mathcal{S})=\langle\operatorname{Sb}(X),T_{i}^{*},R_{ij}^{*}\rangle,$ (29)
where for any binary relation $P\subseteq{}^{2}X$ we define $P^{*}:\operatorname{Sb}(X)\to\operatorname{Sb}(X)$ by
$P^{*}(A)=\{y\in X:\;\langle x,y\rangle\in P\text{ for some }x\in A\}\,.$ (30)
Definition 3.8.
Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}\rangle_{i,j<\alpha}$ be a $cs$-type algebra, such that $\langle A,+,$-$\rangle$ is a Boolean algebra. The ultrafilter frame of $\mathfrak{A}$ is the first-order relational structure
$\mathfrak{Uf}(\mathfrak{A})=\langle{U\!f}(\mathfrak{A}),T_{i},R_{ij}\rangle_{i,j<\alpha},$ (31)
where ${U\!f}(\mathfrak{A})$ is the set of ultrafilters of $\mathfrak{A}$, and $T_{i},R_{ij}\subseteq{}^{2}{U\!f}(\mathfrak{A})$ are defined as
$\langle p,q\rangle\in T_{i}$ $\quad\Longleftrightarrow\quad\{\mathsf{c}_{i}(x):\;x\in p\}\subseteq q,$ (32)
$\langle p,q\rangle\in R_{ij}$ $\quad\Longleftrightarrow\quad\{\mathsf{s}_{ij}(x):\;x\in p\}\subseteq q\,.$ (33)
4. Finitary polyadic algebras and substitution-cylindric algebras
In this section, we introduce the classes $\mathsf{FPA}_{\alpha}$ and $\mathsf{SCA}_{\alpha}$ of algebras defined by finite equational systems.
Definition 4.1 (Cf. Definition 1 in [22]).
By a finitary polyadic algebra of dimension $\alpha$, $\mathsf{FPA}_{\alpha}$, we mean an algebra $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij},\mathsf{p}_{ij}\rangle_{i,j<\alpha}$ in which equations (F0-F9) below are valid for every $i,j,k<\alpha$.
  1. (F0)
    $\langle A,+,$-$\rangle$ is a Boolean algebra [^8] The derived Boolean operations $\cdot$, 0 etc. are taken as defined in the usual way., $\mathsf{s}_{ii}(x)=\mathsf{p}_{ii}(x)=x$, $\mathsf{p}_{ij}(x)=\mathsf{p}_{ji}(x)$.
  2. (F1)
    $x\leq\mathsf{c}_{i}x$
  3. (F2)
    $\mathsf{c}_{i}(x+y)=\mathsf{c}_{i}x+\mathsf{c}_{i}y$
  4. (F3)
    $\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x$
  5. (F4)
    $\mathsf{c}_{i}\mathsf{s}_{ij}x=\mathsf{s}_{ij}x$ if $i\neq j$
  6. (F5)
    $\mathsf{s}_{ij}\mathsf{c}_{k}x=\mathsf{c}_{k}\mathsf{s}_{ij}x$ if $k\notin\{i,j\}$
  7. (F6)
    $\mathsf{s}_{ij}$ and $\mathsf{p}_{ij}$ are Boolean endomorphisms
  8. (F7)
    $\mathsf{p}_{ij}\mathsf{p}_{ij}x=x$
  9. (F8)
    $\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x$ if $i,j,k$ are all distinct
  10. (F9)
    $\mathsf{p}_{ij}\mathsf{s}_{ji}x=\mathsf{s}_{ij}x$
It is routine to check that full finitary polyadic set algebras (18) satisfy 8-(F9), so $\mathsf{RA}_{\alpha}^{csp}\subseteq\mathsf{FPA}_{\alpha}$.
Remark 4.2.
The identities 8-(F9) are positive equations. It is only (F6) which needs a comment. The equation $\mathsf{s}_{ij}(x)=\mathsf{s}_{ij}(-x)$ is part of what it means that $\mathsf{s}_{ij}$ is a Boolean endomorphism, and it is visible that the negation symbol is used here. This equation, as noted in [22, p.561], can be replaced by the ones expressing that $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$. This is because if $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$, then
$\mathsf{s}_{ij}(x)\cdot\mathsf{s}_{ij}(-x)$ $=\mathsf{s}_{ij}(x\cdot-x)=\mathsf{s}_{ij}(0)=0,$ (34)
$\mathsf{s}_{ij}(x)+\mathsf{s}_{ij}(-x)$ $=\mathsf{s}_{ij}(x+-x)=\mathsf{s}_{ij}(1)=1,$ (35)
and these imply $\mathsf{s}_{ij}(-x)=-\mathsf{s}_{ij}(x)$. The same comment applies to the case of $\mathsf{p}_{ij}$. Hence, $\mathsf{FPA}_{\alpha}$ is a canonical variety.
In Section 6, we will use the following consequences of the $\mathsf{FPA}_{\alpha}$-axioms.
Lemma 4.3.
The following identities hold in $\mathsf{FPA}_{\alpha}$.
  1. (P1)
    $\mathsf{p}_{ij}\mathsf{c}_{i}x=\mathsf{s}_{ji}\mathsf{c}_{i}x$
  2. (P2)
    $\mathsf{p}_{ij}\mathsf{c}_{j}x=\mathsf{s}_{ij}\mathsf{c}_{j}x$
  3. (P3)
    $\mathsf{p}_{ij}\mathsf{c}_{k}x=\mathsf{s}_{ik}\mathsf{s}_{ki}\mathsf{s}_{ji}\mathsf{c}_{k}x$ if $k\notin\{i,j\}$
  4. (P4)
    $\mathsf{p}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{j}\mathsf{p}_{ij}x$ for any $i,j<\alpha$
  5. (P5)
    $\mathsf{p}_{ij}\mathsf{c}_{m}x=\mathsf{c}_{m}\mathsf{p}_{ij}x$ if $i,j,m<\alpha$ are distinct.
Proof . Here are the derivations.
[…]
(P4) and (P5) are proved as Lemma 1.6 in [22]. [^9] We note that there are two typos in the statement of Lemma 1.6 in [22]. Namely, (i) writes $\mathsf{p}_{ij}\mathsf{c}_{i}=\mathsf{c}_{i}\mathsf{p}_{ij}$, but the correct version is our (P4). In (iii), the correct assumption is that $\sigma(i)=i$ and for $j\neq i$, $\sigma(j)\neq i$. However, the proof of (the typo-free version of) Lemma 1.6 in [22] is correct, and throughout that paper the lemma is used with the correct statement. $\blacksquare$
Recall that permutations $\sigma:\alpha\to\alpha$ are generated by the transpositions. We let $\mathsf{p}_{[i,j]}=\mathsf{p}_{ij}$ and $\mathsf{p}_{\sigma\circ\tau}=\mathsf{p}_{\sigma}\circ\mathsf{p}_{\tau}$. This way to every permutation $\sigma$ we assign a corresponding polyadic term $\mathsf{p}_{\sigma}$. [22, Claim 1.2] shows that different decompositions of the same permutation into products of transpositions yield identical terms (in $\mathsf{FPA}_{\alpha}$). For a precise definition we refer to [22, p.554]. This leads to the important lemma below that we make use of many times in Section 6.
Lemma 4.4.
The following identity is a consequence of the $\mathsf{FPA}_{\alpha}$ axioms for any permutation $\tau:\alpha\to\alpha$ and $i<\alpha$.
  1. (P6)
    $\mathsf{p}_{\tau}\mathsf{c}_{i}x=\mathsf{c}_{\tau(i)}\mathsf{p}_{\tau}x$
Proof . Apply Lemma 4.3 iteratively. $\blacksquare$
Next we turn to algebras without the $\mathsf{p}_{ij}$ operations. The class $\mathsf{SCA}_{\alpha}$ of algebras is axiomatized by the finitary polyadic algebra axioms of $\mathsf{FPA}_{\alpha}$ without the $\mathsf{p}_{ij}$’s, that is, when the $\mathsf{FPA}_{\alpha}$ axioms speaking about $\mathsf{p}_{ij}$ are deleted. We will see that these algebras are obtained from cylindric algebras by replacing the diagonal constants $\mathsf{d}_{ij}$ with the derived operations $\mathsf{s}_{ij}$. The elements of $\mathsf{SCA}_{\alpha}$ will be called substitution-cylindric algebras.
Definition 4.5 (Cf. Sax on p.559 in [22]).
The class $\mathsf{SCA}_{\alpha}$ consists of the algebras $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}\rangle_{i,j<\alpha}$ in which the equations below are valid for every $i,j,k<\alpha$.
  1. (F0)
    $\langle A,+,$-$\rangle$ is a Boolean algebra, $\mathsf{s}_{ii}(x)=(x)$
  2. (F1)
    $x\leq\mathsf{c}_{i}x$
  3. (F2)
    $\mathsf{c}_{i}(x+y)=\mathsf{c}_{i}x+\mathsf{c}_{i}y$
  4. (F3)
    $\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x$
  5. (F4)
    $\mathsf{c}_{i}\mathsf{s}_{ij}x=\mathsf{s}_{ij}x$ if $i\neq j$
  6. (F5)
    $\mathsf{s}_{ij}\mathsf{c}_{k}x=\mathsf{c}_{k}\mathsf{s}_{ij}x$ if $k\notin\{i,j\}$
  7. (F6)
    $\mathsf{s}_{ij}$ are Boolean endomorphisms
We deliberately use the same labels as in Definition 4.1, even if (F0) and (F6) here are not exactly the same as (F0) and (F6) in Definition 4.1 as here the $\mathsf{p}_{ij}$’s are missing. We believe that this slight abuse of notation will not cause confusion.
Clearly, Remark 4.2 applies to the $\mathsf{SCA}_{\alpha}$ axioms, yielding that these identities are positive, and $\mathsf{SCA}_{\alpha}$ is a canonical variety. Also, $\mathsf{RA}_{\alpha}^{cs}\subseteq\mathsf{SCA}_{\alpha}$.
In the upcoming proofs we will extensively use that diagonal constants $\mathsf{d}_{ij}$ can be adjoined to complete $\mathsf{SCA}_{\alpha}$ algebras so that the cylindric axioms hold. First we recall the definition of a cylindric algebra.
Definition 4.6 (Definition 1.1.1 in [12]).
The class $\mathsf{CA}_{\alpha}$ of cylindric algebras consists of the algebras $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{d}_{ij}\rangle_{i,j<\alpha}$ in which the equations below are valid for every $i,j,k<\alpha$.
  1. (C0)
    $\langle A,+,$-$\rangle$ is a Boolean algebra
  2. (C1)
    $\mathsf{c}_{i}0=0$
  3. (C2)
    $x\leq\mathsf{c}_{i}x$
  4. (C3)
    $\mathsf{c}_{i}(x\cdot\mathsf{c}_{i}y)=\mathsf{c}_{i}x\cdot\mathsf{c}_{i}y$
  5. (C4)
    $\mathsf{c}_{i}\mathsf{c}_{j}x=\mathsf{c}_{j}\mathsf{c}_{i}x$
  6. (C5)
    $\mathsf{d}_{ii}=1$
  7. (C6)
    $\mathsf{d}_{ij}=\mathsf{c}_{k}(\mathsf{d}_{ik}\cdot\mathsf{d}_{kj})$ if $k\neq i,j$
  8. (C7)
    $\mathsf{c}_{i}(\mathsf{d}_{ij}\cdot x)\cdot(\mathsf{d}_{ij}\cdot-x)=0$ if $i\neq j$
Let $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{d}_{ij}\rangle_{i,j<\alpha}\in\mathsf{CA}_{\alpha}$. The substitution-cylindric reduct $\operatorname{Rd}_{sc}\mathfrak{A}$ of $\mathfrak{A}$ is $\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ where $\mathsf{s}_{ij}^{\mathfrak{A}}:A\to A$ is the function defined by $\mathsf{s}_{ij}^{\mathfrak{A}}(x)=\mathsf{c}_{i}(\mathsf{d}_{ij}\cdot x)$ when $i\neq j$ and $\mathsf{s}_{ii}^{\mathfrak{A}}(x)=x$ for $x\in A$, and $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}=\{\operatorname{Rd}_{sc}\mathfrak{A}:\mathfrak{A}\in\mathsf{CA}_{\alpha}\}$. The following theorem says that the variety $\mathsf{SCA}_{\alpha}$ is closely related to the substitution-cylindrification reducts of cylindric algebras.
Theorem 4.7.
Let $3\leq\alpha$ be finite. Then (i)-(ii) below hold.
  1. (i)
    $\mathfrak{A}\in\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$ for every complete $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$.
  2. (ii)
    $\mathsf{SCA}_{\alpha}=\mathbf{S}\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$.
Proof . The proof of (i) relies on results in Pinter [19] and Sain-Thompson [22]. First we prove that (F0)-(F6) imply substitution axioms (S1)-(S6) in [19]. Indeed, Lemma 1.5(iii) and Proposition 6 in [22] prove that the equations
  1. (A1)
    $\mathsf{s}_{ij}\mathsf{s}_{ki}x=\mathsf{s}_{ij}\mathsf{s}_{kj}x$
  2. (A2)
    $\mathsf{s}_{ij}\mathsf{s}_{ik}x=\mathsf{s}_{ik}x$, if $i\neq k$
  3. (A3)
    $\mathsf{s}_{ij}\mathsf{s}_{kl}x=\mathsf{s}_{kl}\mathsf{s}_{ij}x$ if $i\notin\{k,l\}$, $k\neq j$
hold in $\mathsf{SCA}_{\alpha}$ and (S1)-(S6) are (F6), (F0), (A1), (A2), and (A3). Also, [22] proves
$\mathsf{c}_{i}x$ is the least element of $\{y\in\operatorname{ran}\mathsf{s}_{ij}\}$
for $i\neq j$ as (4) on p.515, this is condition ($\pi 1$) + (B) in [19]. Let us define $\mathsf{d}_{ij}^{*}$, for $i\neq j$, as the infimum of $\{y:\mathsf{s}_{ij}y=1\}$ in $\mathfrak{A}$, this infimum exists because $\mathfrak{A}$ is complete. Then (32) on p.561 in [22] proves that $\mathsf{s}_{ij}\mathsf{d}_{ij}^{*}=1$, so
$\mathsf{d}_{ij}^{*}$ is the least element of $\{y:\mathsf{s}_{ij}y=1\}$
[…]
To prove (ii), notice first that $\mathsf{SCA}_{\alpha}\subseteq\mathbf{S}\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$ follows from (i) together with $\mathsf{SCA}_{\alpha}$ being a canonical variety. For the other direction, we have to prove that (F0)-(F6) hold in $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$. Now, (F0) is (C0)+[12, 1.5.1], (F1) is (C2), and (F2)-(F6) are [12, 1.2.6(ii), 1.5.8(i), 1.5.9(ii), 1.5.8(ii), 1.5.3], respectively. $\blacksquare$
5. $\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}^{c}_{\alpha}$
In this section, we prove our first main result that states that a substitution-cylindric algebra of finite dimension $\geq 3$ is representable iff its substitution-free reduct is representable. A corollary to this result is that substitutions over the Booleans together with cylindrifications are finitely axiomatizable (by the axioms of substitution-cylindric algebras). In short: $\mathsf{RA}_{\alpha}^{cs}=\mathsf{RA}_{\alpha}^{c}+\mathsf{SCA}_{\alpha}$.
Theorem 5.1.
For finite $\alpha>2$, an $\alpha$-dimensional algebra $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$ is representable if and only if its substitution-free reduct is representable.
Proof . We show first that it is enough to concentrate on set algebras, i.e., it is enough to prove that if the substitution-free reduct of an $\mathsf{SCA}_{\alpha}$ is a set algebra, then the algebra itself is isomorphic to a set algebra.
Lemma 5.2.
To prove Theorem 5.1 it is enough to show that the statement
if $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$, and $\operatorname{Rd}_{c}\mathfrak{A}\in\mathsf{RA}^{c}_{\alpha}$, then $\mathfrak{A}\in\mathsf{RA}^{cs}_{\alpha}$ ($\sharp$)
holds.
Proof of Lemma 5.2. Assume that ($\sharp$5.2) holds. Take an arbitrary $\mathfrak{A}=\langle A$, +,+,-, $\mathsf{c}_{i}$, $\mathsf{s}_{ij}$ $\rangle_{i,j<\alpha}$ $\in\mathsf{SCA}_{\alpha}$, and assume that $\operatorname{Rd}_{c}\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i}\rangle_{i<\alpha}$ is representable, say
$f:\langle A,+,-,\mathsf{c}_{i}\rangle_{i<\alpha}\quad\hookrightarrow\quad\Pi_{k\in K}\langle\operatorname{Sb}({}^{\alpha}U_{k}),\cup,\smallsetminus,\operatorname{C}_{i}\rangle_{i<\alpha}$ (36)
holds for some $f$ and $U_{k}$, $k\in K$. We have to show that $\mathfrak{A}$ is representable, as well.
Let $k\in K$, denote by $\pi_{k}$ the projection function, and let
$g:\operatorname{Rd}_{c}\mathfrak{A}\rightarrow\langle\operatorname{Sb}({}^{\alpha}U_{k}),\cup,\smallsetminus,\operatorname{C}_{i},\rangle_{i<\alpha}$ (37)
be the composition of $f$ with $\pi_{k}$. Let $\mathfrak{B}_{k}$ be the $g$-image of $\mathfrak{A}$ and let $i,j<\alpha$. We show that (F0)-(F6) (in Definition 4.5 of $\mathsf{SCA}_{\alpha}$) imply that the $g$-image of the function $\mathsf{s}_{ij}$ on $A$ is a function on $B_{k}$. To this end, we have to show that if $g(a)=g(b)$ then $g(\mathsf{s}_{ij}a)=g(\mathsf{s}_{ij}b)$, for any $a,b\in A$. Note that $g$ is a Boolean homomorphism that preserves also the cylindrifications $\mathsf{c}_{i}$. Now, $g(a)=g(b)$ iff, by $g$ being a Boolean homomorphism, $g(a\oplus b)=0$ where $\oplus$ denotes the Boolean symmetric difference. Similarly, $g(\mathsf{s}_{ij}a)=g(\mathsf{s}_{ij}b)$ iff $g(s_{ij}a\oplus\mathsf{s}_{ij}b)=g(\mathsf{s}_{ij}(a\oplus b))=0$, by (F6). Therefore, it is enough to show that $g(\mathsf{s}_{ij}x)=0$ whenever $g(x)=0$. Indeed,
$g(x)=0$ $\quad\Longrightarrow\quad g(\mathsf{c}_{i}x)=\mathsf{c}_{i}g(x)=0$ (by $g$ a $\mathsf{c}_{i}$-homomorphism + $\operatorname{C}_{i}\emptyset=\emptyset$)
$\quad\Longrightarrow\quad g(\mathsf{s}_{ij}\mathsf{c}_{i}x)=0$ (by (F3))
$\quad\Longrightarrow\quad g(\mathsf{s}_{ij}x)\leq g(\mathsf{s}_{ij}\mathsf{c}_{i}x)=0$ (by (F1)+(F6)+ $g$ homomorphism)
$\quad\Longrightarrow\quad g(\mathsf{s}_{ij}x)=0$
We have seen that the $g$-images of the (abstract) substitution operations of $\mathfrak{A}$ are functions on $\mathfrak{B}_{k}$. Let $\mathsf{s}_{ij}^{+}$ denote these projected functions and let $\mathfrak{B}_{k}^{+}=\langle\mathfrak{B}_{k},\mathsf{s}_{ij}^{+}\rangle_{i,j<\alpha}$, for all $k\in K$. Then $g:\mathfrak{A}\to\mathfrak{B}_{k}^{+}$ and so $\mathfrak{B}_{k}^{+}\in\mathsf{SCA}_{\alpha}$ by $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$. Also $\operatorname{Rd}_{c}\mathfrak{B}_{k}^{+}=\mathfrak{B}_{k}\in\mathsf{RA}_{\alpha}^{c}$ by its construction. Thus, $\mathfrak{B}_{k}^{+}\in\mathsf{RA}_{\alpha}^{cs}$, by ($\sharp$5.2) and so $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ since $f:\mathfrak{A}\rightarrowtail\Pi_{k\in K}\mathfrak{B}_{k}^{+}$. That is, $\mathfrak{A}$ is representable. $\blacksquare$
We have seen that it is enough to prove ($\sharp$5.2). To do so, take $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$ such that $\operatorname{Rd}_{c}\mathfrak{A}\in\mathsf{RA}^{c}_{\alpha}$. Then, for some base set $U$ and $A\subseteq\operatorname{Sb}({}^{\alpha}U)$ we can assume, by possibly taking an isomorphic copy of $\mathfrak{A}$, that
$\operatorname{Rd}_{c}\mathfrak{A}=\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U}\rangle_{i<\alpha}.$ (38)
Putting back the substitution operation, we get
$\mathfrak{A}=\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathsf{s}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha},$ (39)
where the operations $\mathsf{s}_{ij}^{\mathfrak{A}}$ are abstract, not yet set-represented operations that we want to represent by the operations $\mathop{S}\nolimits_{ij}$ over sets.
Lemma 5.3.
We can assume that $\mathfrak{A}$ is complete and atomic, suprema are unions and that the $\mathsf{s}_{ij}^{\mathfrak{A}}$’s are completely additive. In short: $\mathfrak{A}$ is a union-complete BAO.
Proof of Lemma 5.3. By Jónsson–Tarski [15] (cf. [4, Thm.2.8.9], [12, 2.7.13]) every $\mathsf{BAO}$, in particular $\mathfrak{A}$, can be embedded into its canonical embedding algebra $\mathfrak{Em}(\mathfrak{A})$. The canonical embedding algebra is union-complete.
Positive equations are preserved by taking the canonical embedding algebra and thus, by Remark 4.2, $\mathfrak{Em}(\mathfrak{A})\in\mathsf{SCA}_{\alpha}$. It only remained to show that $\operatorname{Rd}_{c}\mathfrak{Em}(\mathfrak{A})\in\mathsf{RA}_{\alpha}^{c}$. But this follows from Theorem 3.5, and from $\operatorname{Rd}_{c}\mathfrak{Em}(\mathfrak{A})=\mathfrak{Em}(\operatorname{Rd}_{c}(\mathfrak{A}))$ (straightforward by definition). Finally, if $\mathfrak{Em}(\mathfrak{A})$ is representable, then so is $\mathfrak{A}$ by $\mathfrak{A}$ being a subalgebra of $\mathfrak{Em}(\mathfrak{A})$. $\blacksquare$
From now on, we assume that $\mathfrak{A}$ is as in Lemma 5.3. By Theorem 4.7(i) then $\mathfrak{A}\in\operatorname{Rd}_{sc}\mathfrak{B}$ for some $\mathfrak{B}\in\mathsf{CA}_{\alpha}$, since $\mathfrak{A}\in\mathsf{SCA}_{\alpha}$ is complete. By the definition of $\operatorname{Rd}_{sc}\mathfrak{B}$ this means that there are $\mathsf{d}_{ij}^{\mathfrak{A}}\in A$ for $i,j<\alpha$ such that $\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is a cylindric algebra, where
$\mathsf{s}_{ij}^{\mathfrak{A}}x=\begin{cases}x&\mbox{ if }i=j,\\ \operatorname{C}_{i}^{U}(x\cap\mathsf{d}_{ij}^{\mathfrak{A}})&\mbox{ if }i\neq j\end{cases}$ (40)
holds.
We now change the representation of $\mathfrak{A}$ in order to arrange that these “abstract” diagonal constants $\mathsf{d}_{ij}^{\mathfrak{A}}$ be represented close to the set diagonal constants $\operatorname{D}_{ij}^{U}$. We cannot achieve $\mathsf{d}_{ij}=\operatorname{D}_{ij}$ because the diagonals $\operatorname{D}_{ij}$ are not finitely axiomatizable over $\mathsf{RA}^{cs}$, see [1, Theorem 6]. We will be able to achieve $\operatorname{D}_{ij}\subseteq\mathsf{d}_{ij}$ and this will be enough for representing the $\mathsf{s}_{ij}$ well. The idea for achieving $\operatorname{D}_{ij}\subseteq\mathsf{d}_{ij}$ is that first we “blow up” each element of $U$ and then we will select a bijection below each $\mathsf{d}_{ij}$ along which we will “rearrange” the sequences so that we get $\operatorname{D}_{ij}\subseteq\mathsf{d}_{ij}$.
We now change the representaton by blowing up each element of $U$. Let $W$ be an infinite set such that $|W|>|U|$. The new base set of the representation will be $V=U\times W$. For $s\in{}^{\alpha}U$ and $w\in{}^{\alpha}W$ let $s\times w\in{}^{\alpha}V$ be defined as
$s\times w=\langle(s_{i},w_{i}):i<\alpha\rangle.$
For each sequence $s\in{}^{\alpha}U$ let us write
$\hat{s}=\{s\times w:\;w\in{}^{\alpha}W\},$ (41)
and for $X\in\operatorname{Sb}({}^{\alpha}U)$ let
$F(X)=\bigcup\big\{\hat{s}:\;s\in X\big\}\in\operatorname{Sb}({}^{\alpha}V).$ (42)
It is straightforward to check that $F$ is an embedding
$\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U}\rangle_{i<\alpha}\quad\hookrightarrow\quad\langle\operatorname{Sb}({}^{\alpha}V),\cup,\smallsetminus,\operatorname{C}_{i}^{V}\rangle_{i<\alpha}$ (43)
that takes infinite unions to unions, i.e., $F(\bigcup X)=\bigcup\{F(x):x\in X\}$ for all $X\subseteq A$. Consequently, there is a subset $A^{\prime}\subseteq\operatorname{Sb}({}^{\alpha}V)$ such that $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle$ is isomorphic to an algebra
$\mathfrak{A}^{\prime}=\langle A^{\prime},\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathsf{s}_{ij}^{\prime},\mathsf{d}_{ij}^{\prime}\rangle_{i<\alpha},$ (44)
where $\mathsf{s}_{ij}^{\prime}$ and $\mathsf{d}_{ij}^{\prime}$ are the isomorphic images of $\mathsf{s}_{ij}^{\mathfrak{A}}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, respectively. Therefore, $\mathfrak{A}^{\prime}$ is a union-complete BAO, with the corresponding reducts being a cylindric algebra and a substitution-cylindric algebra, with $\mathsf{s}_{ij}^{\prime}$ defined as in (40).
We now make preparations for changing the representation once more in order to achieve $\operatorname{D}_{ij}\subseteq\mathsf{d}_{ij}$. Elements of $\operatorname{Sb}({}^{\alpha}V)$ are $\alpha$-ary relations. For such a relation $R\subseteq{}^{\alpha}V$ and $i<\alpha$ we define
$\operatorname{dom}_{i}(R)=\{s(i):\;s\in R\}\ \subseteq\ V.$ (45)
In what follows, $At$ is the set of atoms of $\mathfrak{A}^{\prime}$.
Lemma 5.4.
Let $i\neq j<\alpha$, and let $a,b\subseteq\mathsf{d}_{ij}^{\prime}$ be two atoms. Then either
$\operatorname{dom}_{i}(a)=\operatorname{dom}_{i}(b),\quad\text{ and }\quad\operatorname{dom}_{j}(a)=\operatorname{dom}_{j}(b),$ (46)
or
$\operatorname{dom}_{i}(a)\cap\operatorname{dom}_{i}(b)=\emptyset,\quad\text{ and }\quad\operatorname{dom}_{j}(a)\cap\operatorname{dom}_{j}(b)=\emptyset.$ (47)
Proof of Lemma 5.4. Let $a,b$ be distinct atoms below $\mathsf{d}_{ij}^{\prime}$, and set $\Gamma=\alpha-\{i,j\}$. Below, $\operatorname{C}_{\Gamma}^{V}(x)$ denotes $\operatorname{C}_{i1}^{V}\dots\operatorname{C}_{ig}^{V}(x)$ where $\Gamma=\{i1,\dots,ig\}$. This notation makes sense by the properties of $\operatorname{C}_{i}$. By [12, 1.3.3], $\operatorname{C}_{\Gamma}^{V}(\mathsf{d}_{ij}^{\prime})=\mathsf{d}_{ij}^{\prime}$, and by [12, 1.10.3], $\operatorname{C}_{\Gamma}^{V}(a)$ is an atom of the algebra of $\Gamma$-closed elements. In particular, $\operatorname{C}_{\Gamma}^{V}(a)$ and $\operatorname{C}_{\Gamma}^{V}(b)$ are below $\mathsf{d}_{ij}^{\prime}$, and are either equal or disjoint. By definition of $\operatorname{C}_{\Gamma}^{V}$, we have
$\operatorname{dom}_{i}(a)=\operatorname{dom}_{i}(\operatorname{C}_{\Gamma}^{V}(a)),\quad\operatorname{dom}_{j}(a)=\operatorname{dom}_{j}(\operatorname{C}_{\Gamma}^{V}(a)),$ (48)
$\operatorname{dom}_{i}(b)=\operatorname{dom}_{i}(\operatorname{C}_{\Gamma}^{V}(b)),\quad\operatorname{dom}_{j}(b)=\operatorname{dom}_{j}(\operatorname{C}_{\Gamma}^{V}(b))\,.$ (49)
If $\operatorname{C}_{\Gamma}^{V}(a)=\operatorname{C}_{\Gamma}^{V}(b)$, then (46) follows. In the other case, assume that
$\operatorname{C}_{\Gamma}^{V}(a)\cap\operatorname{C}_{\Gamma}^{V}(b)=\emptyset\,.$ (50)
We show (47). By way of contradiction, suppose there is $v\in\operatorname{dom}_{i}(a)\cap\operatorname{dom}_{i}(b)$. Then there are $x,y\in{}^{\alpha}V$ such that $x\in a$, $y\in b$, and $x(i)=y(i)=v$. Let $z=x(j/y(j))$. Then $z\in\operatorname{C}_{\Gamma}^{V}(b)$, because $z$ and $y$ can differ only at their $\Gamma$-coordinates. Also, $z\in\operatorname{C}_{j}^{V}(a)$, because $x$ and $z$ may differ only at the $j$th coordinate. But then
$z$ $\in\mathsf{d}_{ij}^{\prime}\cap-\operatorname{C}_{\Gamma}^{V}(a)$ (by (50) and $z\in\operatorname{C}_{\Gamma}^{V}(b)\subseteq\mathsf{d}_{ij}^{\prime}$)
$z$ $\in\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime}\cap\operatorname{C}_{\Gamma}^{V}(a))$ (by $\operatorname{C}_{\Gamma}^{V}(a)\subseteq\mathsf{d}_{ij}^{\prime}$, monotonicity of $\operatorname{C}_{\Gamma}^{V}$ and $z\in\operatorname{C}_{j}^{V}(a)$)
Using monotonicity of $\operatorname{C}_{j}^{V}$ we end up at
$z\in\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime}\cap\operatorname{C}_{\Gamma}^{V}(a))\cap\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime}\cap-\operatorname{C}_{\Gamma}^{V}(a)),$ (51)
which contradicts cylindric axiom (C7) (see Definition 4.6):
$\mathsf{CA}_{\alpha}\vDash\mathsf{c}_{j}(\mathsf{d}_{ij}\cdot x)\cdot\mathsf{c}_{j}(\mathsf{d}_{ij}\cdot-x)=0.$ (52)
We proved that if (50) holds, then $\operatorname{dom}_{i}(a)\cap\operatorname{dom}_{i}(b)=\emptyset$. Exchanging $i$ and $j$ in the argument yields $\operatorname{dom}_{j}(a)\cap\operatorname{dom}_{j}(b)=\emptyset$. $\blacksquare$
Lemma 5.5.
For all $i,j<\alpha$, $\operatorname{dom}_{i}(\mathsf{d}_{ij}^{\prime})=\operatorname{dom}_{j}(\mathsf{d}_{ij}^{\prime})=V$.
Proof of Lemma 5.5. By the cylindric axioms (see [12, 1.3.2]),
$\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime})={}^{\alpha}V\,.$ (53)
Now, if for some $v\in V$, $v\notin\operatorname{dom}_{i}(\mathsf{d}_{ij}^{\prime})$, then $v\notin\operatorname{dom}_{i}\big(\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime})\big)$, as the $i$th coordinates of the sequences in $\mathsf{d}_{ij}^{\prime}$ and $\operatorname{C}_{j}^{V}(\mathsf{d}_{ij}^{\prime})$ are the same. But then (53) would be impossible. $\blacksquare$
Fix $i\neq 0$ and let $a\subseteq\mathsf{d}_{0i}^{\prime}$ be an atom. By construction (of the embedding $F$), if $\langle u,w\rangle\in\operatorname{dom}_{0}(a)$, then $\langle u,v\rangle\in\operatorname{dom}_{0}(a)$ for every $v\in W$, and similarly with $\operatorname{dom}_{i}(a)$. This means that every $\langle u,w\rangle\in\operatorname{dom}_{0}(a)$ has $|W|$-many $a$-neighbours in $\operatorname{dom}_{i}(a)$, and vice-versa. Using transfinite recursion, we construct a bijective function
[…]
To do so, enumerate $\operatorname{dom}_{0}(a)$ and $\operatorname{dom}_{i}(a)$ in $|W|$-type, and starting with the empty list of pairs, in consecutive rounds for each not yet selected element in $\operatorname{dom}_{0}(a)$ pick one of its $a$-neighbours in $\operatorname{dom}_{i}(a)$ that has not yet been selected; and for each not yet selected element in $\operatorname{dom}_{i}(a)$ find a pair among its $a$-neighbours in $\operatorname{dom}_{0}(a)$ that has not yet been picked. This is possible, because each element has $|W|$ possible pairs in $a$, and $|W|>|U|$ is infinite. This transfinite back and forth procedure gives us a pairing of elements (i.e., a bijection), and as the entire $\operatorname{dom}_{0}(a)$ and $\operatorname{dom}_{i}(a)$ were enumerated, $\operatorname{dom}(f_{0i}^{a})=\operatorname{dom}_{0}(a)$, and $\operatorname{ran}(f_{0i}^{a})=\operatorname{dom}_{i}(a)$.
If $a,b\subseteq\mathsf{d}_{0i}^{\prime}$ are distinct atoms, then we have two cases according to Lemma 5.4: either
$\operatorname{dom}_{0}(a)=\operatorname{dom}_{0}(b),\quad\text{ and }\quad\operatorname{dom}_{i}(a)=\operatorname{dom}_{i}(b),$ (55)
or
$\operatorname{dom}_{0}(a)\cap\operatorname{dom}_{0}(b)=\emptyset,\quad\text{ and }\quad\operatorname{dom}_{i}(a)\cap\operatorname{dom}_{i}(b)=\emptyset.$ (56)
We construct $f_{0i}^{a}$ and $f_{0i}^{b}$ in such a way that in case (55) we pick the same bijections: If (55) holds for the atoms $a$ and $b$, then we choose $f_{0i}^{a}=f_{0i}^{b}$.
Define
$f_{0i}=\bigcup\{f_{0i}^{a}:\;a\in At,\ a\subseteq\mathsf{d}_{0i}^{\prime}\},\quad f_{00}=\operatorname{id}.$ (57)
As the union of atoms below $\mathsf{d}_{0i}^{\prime}$ is $\mathsf{d}_{0i}^{\prime}$ itself (the algebra is a union-complete BAO), $f_{0i}$ is a relation with $\operatorname{dom}(f_{0i})=\operatorname{dom}_{0}(\mathsf{d}_{0i}^{\prime})$, and $\operatorname{ran}(f_{0i})=\operatorname{dom}_{i}(\mathsf{d}_{0i}^{\prime})$.
Lemma 5.6.
$f_{0i}$ is a bijective function from $V$ to $V$.
Proof of Lemma 5.6. Combining $\operatorname{dom}(f_{0i})=\operatorname{dom}_{0}(\mathsf{d}_{0i}^{\prime})$, and $\operatorname{ran}(f_{0i})=\operatorname{dom}_{i}(\mathsf{d}_{0i}^{\prime})$ with Lemma 5.5, we get $\operatorname{dom}(f_{0i})=V$ and $\operatorname{ran}(f_{0i})=V$. By Lemma 5.4 and Lemma 5.5, the sets
$\big\{\operatorname{dom}(f_{0i}^{a}):\;a\in At,\ a\subseteq\mathsf{d}_{0i}^{\prime}\big\},\text{ and }\big\{\operatorname{ran}(f_{0i}^{a}):\;a\in At,\ a\subseteq\mathsf{d}_{0i}^{\prime}\big\}$ (58)
are partitions of $V$. Lemma 5.4 also implies that it is not possible to have atoms $a,b\subseteq\mathsf{d}_{ij}^{\prime}$ such that $\operatorname{dom}_{0}(a)=\operatorname{dom}_{0}(b)$ but $\operatorname{dom}_{i}(a)\neq\operatorname{dom}_{i}(b)$; or $\operatorname{dom}_{0}(a)\neq\operatorname{dom}_{0}(b)$ but $\operatorname{dom}_{i}(a)=\operatorname{dom}_{i}(b)$. Since in the case $\operatorname{dom}_{0}(a)=\operatorname{dom}_{0}(b)$ and $\operatorname{dom}_{i}(a)=\operatorname{dom}_{i}(b)$ we chose $f_{0i}^{a}=f_{0i}^{b}$, it follows that $f_{0i}$ is the union of bijections having pairwise disjoint domains and pairwise disjoint ranges. Putting all this together, we get that $f_{0i}$ is a bijection. $\blacksquare$
In the next step, we extend $f_{0i}$ to a bijection $f:{}^{\alpha}V\to{}^{\alpha}V$ by letting for every $s\in{}^{\alpha}V$,
$f(s)=\langle f_{0i}^{-1}\big(s(i)\big):\;i<\alpha\rangle\in{}^{\alpha}V\,.$ (59)
This $f$ induces the mapping
$G(X)=\big\{x\in{}^{\alpha}V:\;f(x)\in X\big\}.$ (60)
Lemma 5.7.
The mapping $G$ is an isomorphism
[…]
where $A^{\prime\prime}$, $\mathsf{s}_{ij}^{\prime\prime}$, and $\mathsf{d}_{ij}^{\prime\prime}$ are the $G$-images of $\mathsf{s}_{ij}^{\prime}$ and $\mathsf{d}_{ij}^{\prime}$, respectively. Further, $G$ takes infinite unions to unions.
The content of the lemma is that the set theoretic operations $\operatorname{C}_{i}^{V}$ are kept fixed under the isomorphism $G$. This lemma essentially follows from [12, Theorem 5.1.34], nevertheless we provide the proof.
Proof of Lemma 5.7. The case of Boolean operations is straightforward, and thus we only need to check $G(\operatorname{C}_{i}^{V}(R))=\operatorname{C}_{i}^{V}(R)$ for $R\subseteq{}^{\alpha}V$. Unfolding the definitions yields
$s\in\operatorname{C}_{i}^{V}(G(R))$ $\Leftrightarrow(\exists u\in V)f(s(i/u))\in R,$ (64)
$s\in G(\operatorname{C}_{i}^{V}(R))$ $\Leftrightarrow(\exists u\in V)f(s)(i/u)\in R.$ (65)
Now, by definition of $f$ we have
$f\big(s(i/u)\big)(j)=f_{0j}^{-1}\big(s(i/u)(j)\big)=\begin{cases}f_{0j}^{-1}\big(s(j)\big)&\text{ if }i\neq j,\\ f_{0j}^{-1}\big(u\big)&\text{ otherwise.}\end{cases}$ (66)
On the other hand,
$\big(f(s)(i/u)\big)(j)=\begin{cases}f(s)(j)=f_{0j}^{-1}\big(s(j)\big)&\text{ if }i\neq j,\\ u&\text{ otherwise.}\end{cases}$ (67)
As each $f_{0j}^{-1}$ is a bijection, it follows that $(\exists u\in V)f(s(i/u))\in R$ holds if and only if $(\exists u\in V)f(s)(i/u)\in R$ does.
Finally, it is routine to check that $G$ takes infinite unions to unions. $\blacksquare$
Lemma 5.8.
For every $i,j<\alpha$ we have $\operatorname{D}_{ij}^{V}\subseteq\mathsf{d}_{ij}^{\prime\prime}$.
Proof of Lemma 5.8. We show first $\operatorname{D}_{0i}^{V}\subseteq\mathsf{d}_{0i}^{\prime\prime}$ for $i<\alpha$. Recall from Lemma 5.6 that $\operatorname{dom}(f_{0i})=\operatorname{ran}(f_{0i})=V$, and by construction
$f_{0i}\times{}^{\alpha-2}V=\{s\in{}^{\alpha}V:s(0)\in\operatorname{dom}(f_{0i}),\ s(i)=f_{0i}(s(0))\}\ \subseteq\ \mathsf{d}_{0i}^{\prime}\,.$ (68)
(By cylindric axioms, $\operatorname{C}_{k}^{V}\mathsf{d}_{0i}^{\prime}=\mathsf{d}_{0i}^{\prime}$ for $k\neq 0,i$, thus, sequences in $\mathsf{d}_{0i}^{\prime}$ can take any values on the coordinates different from 0 and $i$). Assume $s\in{}^{\alpha}V$, $s\in\operatorname{D}_{0i}^{V}$, that is, $s(0)=s(i)$. We need to prove $s\in\mathsf{d}_{0i}^{\prime\prime}$. As $f$ is a bijection, there is a unique $z\in{}^{\alpha}V$ such that $f(z)=s$. By definition of $f$, we have $s(i)=f_{0i}^{-1}\big(z(i)\big)$, in particular, $s(0)=z(0)$ (because $f_{00}=\operatorname{id}$). Combining this with $s(0)=s(i)$, we get $z(0)=f_{0i}^{-1}(z(i))$, that is,
$z(i)=f_{0i}(z(0))\,.$ (69)
But this means that $z\in\mathsf{d}_{0i}^{\prime}$, by (68). Therefore $s\in f(\mathsf{d}_{0i}^{\prime})=\mathsf{d}_{0i}^{\prime\prime}$. We proved $\operatorname{D}_{0i}^{V}\subseteq\mathsf{d}_{0i}^{\prime\prime}$.
The rest of the cases needs axioms of cylindric algebras. $\operatorname{D}_{i0}^{V}\subseteq\mathsf{d}_{i0}^{\prime\prime}$ because $\operatorname{D}_{i0}^{V}=\operatorname{D}_{0i}^{V}$, $\operatorname{D}_{0i}^{V}\subseteq\mathsf{d}_{0i}^{\prime\prime}$, and $\mathsf{CA}_{\alpha}\vDash\mathsf{d}_{0i}=\mathsf{d}_{i0}$ ([12, 1.3.1]), in particular, $\mathsf{d}_{0i}^{\prime\prime}=\mathsf{d}_{i0}^{\prime\prime}$. Finally, let $i,j<\alpha$, $i,j,0$ distinct. Then
$\operatorname{D}_{ij}^{V}=\operatorname{C}_{0}^{V}(\operatorname{D}_{i0}^{V}\cap\operatorname{D}_{0j}^{V})\subseteq\operatorname{C}_{0}^{V}(\mathsf{d}_{i0}^{\prime\prime}\cap\mathsf{d}_{0j}^{\prime\prime})=\mathsf{d}_{ij}^{\prime\prime}$ (70)
by cylindric axiom (C6). $\blacksquare$
So far, we have a cylindric algebra
$\mathfrak{A}^{\prime\prime}=\langle A^{\prime\prime},\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathsf{s}_{ij}^{\prime\prime},\mathsf{d}_{ij}^{\prime\prime}\rangle_{i<\alpha},$ (71)
such that $\operatorname{Rd}_{cs}(\mathfrak{A}^{\prime\prime})\cong\operatorname{Rd}_{cs}(\mathfrak{A})$; and $\mathfrak{A}^{\prime\prime}$ is complete, atomic, suprema are unions, and each $\mathsf{s}_{ij}^{\prime\prime}$ is completely additive satisfying
$\mathsf{s}_{ij}^{\prime\prime}(x)=\begin{cases}x&\mbox{ if }i=j,\\ \operatorname{C}_{i}^{V}(x\cap\mathsf{d}_{ij}^{\prime\prime})&\mbox{ if }i\neq j,\end{cases}$ (72)
and $\operatorname{D}_{ij}^{V}\subseteq\mathsf{d}_{ij}^{\prime\prime}$.
Next we prove that $\mathsf{s}_{ij}^{\prime\prime}$ are the real, set-theoretic operations $\mathop{S}\nolimits_{ij}^{V}$.
Lemma 5.9.
For all $i,j<\alpha$, $\mathsf{s}_{ij}^{\prime\prime}=\mathop{S}\nolimits_{ij}^{V}$.
Proof of Lemma 5.9. The case $i=j$ is straightforward: $\mathsf{s}_{ii}^{\prime\prime}=\operatorname{id}=\mathop{S}\nolimits_{ii}^{V}$. Take $i\neq j<\alpha$. As the algebra is complete, atomic, and $\mathsf{s}_{ij}^{\prime\prime}$ is completely additive, $\mathsf{s}_{ij}^{\prime\prime}$ is determined by its values on atoms. Recall the definitions
$\operatorname{C}_{i}^{V}(a)$ $=\{x\in{}^{\alpha}V:\;x(i/v)\in a\text{ for some }v\in V\},$ (73)
$\mathop{S}\nolimits_{ij}^{V}(a)$ $=\{x\in{}^{\alpha}V:\;x(i/x(j))\in a\}.$ (74)
By (72) we have $\mathsf{s}_{ij}^{\prime\prime}(x)=\operatorname{C}_{i}^{V}(x\cap\mathsf{d}_{ij}^{\prime\prime})$.
For an atom $a\not\subseteq\mathsf{d}_{ij}$ we have $\mathsf{s}_{ij}^{\prime\prime}(a)=\emptyset$ by $a\cap\mathsf{d}_{ij}=\emptyset$, (72) and $\operatorname{C}_{i}^{V}\emptyset=\emptyset$. As $\operatorname{D}_{ij}^{V}\subseteq\mathsf{d}_{ij}^{\prime\prime}$ we also get $\mathop{S}\nolimits_{ij}^{V}(a)=\emptyset$. Therefore, what we need to check is whether
$\mathsf{s}_{ij}^{\prime\prime}(a)=\mathop{S}\nolimits_{ij}^{V}(a)\quad\text{ for }a\subseteq\mathsf{d}_{ij}^{\prime\prime}\text{ atoms}.$ (75)
Now, if $a$ is an atom below $\mathsf{d}_{ij}^{\prime\prime}$, then $\mathsf{s}_{ij}^{\prime\prime}(a)=\operatorname{C}_{i}^{V}(a\cap\mathsf{d}_{ij}^{\prime\prime})=\operatorname{C}_{i}^{V}(a)$, and so we need to prove
$\operatorname{C}_{i}^{V}(a)=\mathop{S}\nolimits_{ij}^{V}(a)\quad\text{ for }a\subseteq\mathsf{d}_{ij}^{\prime\prime}\text{ atoms}.$ (76)
The inclusion $\mathop{S}\nolimits_{ij}^{V}\subseteq\operatorname{C}_{i}^{V}$ is immediate from the definitions. For the other direction, take an arbitrary $s\in\operatorname{C}_{i}^{V}(a)$, and let
$z=s\big(i/s(j)\big)\,.$ (77)
To get $s\in\mathop{S}\nolimits_{ij}^{V}(a)$ we need $z\in a$. Now, $z\in D_{ij}^{V}$ as $z(i)=s(j)=z(j)$. As $\operatorname{D}_{ij}^{V}\subseteq\mathsf{d}_{ij}^{\prime\prime}$, we get $z\in\mathsf{d}_{ij}^{\prime\prime}$.
Suppose, by way of contradiction, that $z\notin a$. Then $z\in-a\cap\mathsf{d}_{ij}^{\prime\prime}$, and thus $s\in\operatorname{C}_{i}^{V}(-a\cap\mathsf{d}_{ij}^{\prime\prime})$. On the other hand, $s\in\operatorname{C}_{i}^{V}(a)$, and $a\subseteq\mathsf{d}_{ij}^{\prime\prime}$, so $s\in\operatorname{C}_{i}^{V}(a\cap\mathsf{d}_{ij}^{\prime\prime})$. It follows that
$s\in\operatorname{C}_{i}^{V}(-a\cap\mathsf{d}_{ij}^{\prime\prime})\cap\operatorname{C}_{i}^{V}(a\cap\mathsf{d}_{ij}^{\prime\prime}).$ (78)
But this contradicts cylindric axiom (C7). Consequently, $z\in a$, and this completes the proof of Lemma 5.9. $\blacksquare$
Summing up, we have shown that
$\langle A^{\prime\prime},\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathsf{s}_{ij}^{\prime\prime}\rangle_{i<\alpha}=\langle A^{\prime\prime},\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathop{S}\nolimits_{ij}^{V}\rangle_{i<\alpha}\in\mathsf{RA}^{cs}_{\alpha}.$ (79)
This implies $\mathfrak{A}\in\mathsf{RA}^{cs}_{\alpha}$, which makes the proof of Theorem 5.1 complete. $\blacksquare$
As a corollary to Theorem 5.1, we can fill in the status of the open line from $c$ to $cs$ in [1, Figure 1]: it should be a dashed line indicating finite axiomatizability.
Theorem 5.10.
$\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$ (by the equations (F0)-(F6) defining $\mathsf{SCA}_{\alpha}$).
Proof . This is an immediate corollary of Lemma 5.2 and $\mathsf{RA}_{\alpha}^{cs}\subseteq\mathsf{SCA}_{\alpha}$. $\blacksquare$
We now prove Theorem 1.2. Recall first that $\mathsf{Rdf}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and $\mathsf{RFPA}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$. Then $\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}=\mathbf{S}\mathbf{SP}\operatorname{Rd}_{cs}\mathsf{RA}_{\alpha}^{csp}=\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$. So Theorem 1.2 states that $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ iff ($\operatorname{Rd}_{c}\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and equations (6)–(9) hold in $\mathfrak{A}$). By Theorem 5.10 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{c}$ and equations (F0)-(F6) hold in $\mathfrak{A}$). Equations (6)–(9) in Theorem 1.2 are the equations in (F0)-(F6) in which the substitutions occur [^11] The indices $i,j,k$ in Theorem 1.2 are required to be distinct while in (F0)–(F6) they are not required to be distinct. It is easy to check that this does not make any difference.; the ones in which the substitutions do not occur hold in $\mathsf{RA}_{\alpha}^{c}$, so they are superfluous. Theorem 1.2 has been proved.
6. $\mathsf{RA}_{\alpha}^{csp}$ is finitely axiomatizable over $\mathsf{RA}^{cs}_{\alpha}$
This section proves our second main result which states that a finitary polyadic algebra of finite dimension is representable iff its permutation-free reduct is representable. A corollary to this result is that permutations over the Booleans together with substitutions and cylindrifications are finitely axiomatizable (by the axioms of finitary polyadic algebras). In short: $\mathsf{RA}_{\alpha}^{csp}=\mathsf{RA}_{\alpha}^{cs}+\mathsf{FPA}_{\alpha}$.
Theorem 6.1.
For finite $\alpha>2$, an $\alpha$-dimensional algebra $\mathfrak{A}\in\mathsf{FPA}_{\alpha}$ is representable if and only if its permutation-free reduct is representable.
Proof . The beginning of the proof follows the proof of Theorem 5.1. Namely, it is enough to prove that if the permutation-free reduct of an $\mathsf{FPA}_{\alpha}$ is isomorphic to a set algebra, then the algebra itself is isomorphic to a set algebra:
if $\mathfrak{A}\in\mathsf{FPA}_{\alpha}$, and $\operatorname{Rd}_{cs}\mathfrak{A}\in\mathsf{RA}^{cs}_{\alpha}$, then $\mathfrak{A}\in\mathsf{RA}^{csp}_{\alpha}$ (80)
The proof of Lemma 5.2 with the obvious modifications works to prove this, too. (In showing $g(x)=0\Rightarrow g(\mathsf{p}_{ij}x)=0$ one uses (F7) in place of (F3).) To prove (80), take $\mathfrak{A}\in\mathsf{FPA}_{\alpha}$ such that $\operatorname{Rd}_{cs}\mathfrak{A}\in\mathsf{RA}^{cs}_{\alpha}$. Then, for some base set $U$ and $A\subseteq\operatorname{Sb}({}^{\alpha}U)$ we can assume, by possibly taking an isomorphic copy of $\mathfrak{A}$, that
$\operatorname{Rd}_{cs}\mathfrak{A}=\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U}\rangle_{i,j<\alpha}.$ (81)
Putting back the permutation operations, we get
$\mathfrak{A}=\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U},\mathsf{p}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha},$ (82)
where the operations $\mathsf{p}_{ij}^{\mathfrak{A}}$ are abstract, not yet set-represented operations that we want to represent by the operations $\mathop{P}\nolimits_{ij}$ over sets. As before, we can assume that $\mathfrak{A}$ is complete and atomic, suprema are unions and that the $\mathsf{p}_{ij}^{\mathfrak{A}}$’s are completely additive (by taking $\mathfrak{Em}(\mathfrak{A})$ in place of $\mathfrak{A}$). As in the analogous step in the proof of Theorem 5.1, we can find elements $\mathsf{d}_{ij}^{\mathfrak{A}}\in A$ for $i,j<\alpha$ such that $\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is a cylindric algebra, and
$\mathop{S}\nolimits_{ij}^{U}x=\begin{cases}x&\mbox{ if }i=j,\\ \operatorname{C}_{i}^{U}(x\cap\mathsf{d}_{ij}^{\mathfrak{A}})&\mbox{ if }i\neq j\end{cases}$ (83)
holds. We also have
$\operatorname{D}_{ij}^{U}\subseteq\mathsf{d}_{ij}^{\mathfrak{A}}$ (84)
since (83) implies $\mathop{S}\nolimits_{ij}^{U}(-\mathsf{d}_{ij}^{\mathfrak{A}})=\emptyset$, so $\operatorname{D}_{ij}^{U}\cap-\mathsf{d}_{ij}^{\mathfrak{A}}=\emptyset$ by the definition of $\mathop{S}\nolimits_{ij}^{U}$.
Lemma 6.2.
$\mathsf{p}_{ij}^{\mathfrak{A}}(\mathsf{d}_{kl}^{\mathfrak{A}})=\mathsf{d}_{\tau(k)\tau(l)}^{\mathfrak{A}}$ where $\tau=[i,j]$.
Proof of Lemma 6.2. For any $i,j<\alpha$, from (83), [12, 1.5.4(i)] and cylindric axiom (C2) we get $\mathop{S}\nolimits_{ij}^{U}(\mathsf{d}_{ij}^{\mathfrak{A}})=1$ and $x\cap\mathsf{d}_{ij}^{\mathfrak{A}}\subseteq\mathop{S}\nolimits_{ij}^{U}(x)$. Sain–Thompson [22] define the class $\mathsf{FPEA}_{\alpha}$ as $\mathsf{FPA}_{\alpha}$ enriched with diagonal elements $\mathsf{d}_{ij}$ satisfying $\mathsf{s}_{ij}\mathsf{d}_{ij}=1$ and $x\cdot\mathsf{d}_{ij}\leq\mathsf{s}_{ij}x$ (see [22, p.544, Definition 2]). We just verified that these two properties hold for $\mathop{S}\nolimits_{ij}^{U}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, therefore the expanded structure $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is an $\mathsf{FPEA}_{\alpha}$. [22] proves that $\mathsf{FPEA}_{\alpha}$ are term equivalent to quasi-polyadic equality algebras $\mathsf{QPEA}_{\alpha}$, and in particular, [22, p.557-558, (E3)] proves that from $\mathsf{FPEA}_{\alpha}$ the identity
$\mathsf{p}_{ij}(\mathsf{d}_{kl})=\mathsf{d}_{\tau(k)\tau(l)}$ (85)
follows. $\blacksquare$
Write $At$ for the atoms of $\mathfrak{A}$, and let $\mathfrak{B}$ be the Boolean algebra generated by the elements $\operatorname{C}_{i}^{U}(a)$ for $i<\alpha$ and $a\in At$:
$\mathfrak{B}=\mathfrak{Sg}^{\mathfrak{Bl}\mathfrak{A}}\big(\{\operatorname{C}_{i}^{U}(a):\;a\in At,i<\alpha\}\big)\,.$ (86)
Lemma 6.3.
$\mathsf{p}_{ij}^{\mathfrak{A}}(x)=\mathop{P}\nolimits_{ij}^{U}(x)$ for every $x\in B$ and $i,j<\alpha$.
Proof of Lemma 6.3. For $i=j$ the statement follows from 8 $\mathsf{p}_{ii}^{\mathfrak{A}}=\operatorname{id}=\mathop{P}\nolimits_{ii}^{U}$. Let $i\neq j$. As $\mathsf{p}_{ij}^{\mathfrak{A}}$ is a Boolean endomorphism by (F6), it is enough to check
$\mathsf{p}_{ij}^{\mathfrak{A}}(\operatorname{C}_{k}^{U}(a))=\mathop{P}\nolimits_{ij}^{U}(\operatorname{C}_{k}^{U}(a))$ (87)
for $a\in At$ and $k<\alpha$. We have three cases depending on whether $k=i$, $k=j$, or $k\neq i,j$. The identities (P1), (P2), and (P3) can be used for these cases. For example, if $k=i$, then
[…]
where the last equality is a property of the set theoretic $\mathop{P}\nolimits_{ij}^{U}$. The rest of the cases are similar. $\blacksquare$
Lemma 6.4.
$\mathsf{p}_{ij}^{\mathfrak{A}}(a)=\mathop{P}\nolimits_{ij}^{U}(a)$ for every atom $a\in At$, $a\subseteq\mathsf{d}^{\mathfrak{A}}_{kl}$, $k\neq l<\alpha$.
Proof of Lemma 6.4. By (F6), $\mathsf{p}_{ij}^{\mathfrak{A}}$ is a Boolean endomorphism, and by (F7) $\mathsf{p}_{ij}^{\mathfrak{A}}$ is invertible. Therefore, $\mathsf{p}_{ij}^{\mathfrak{A}}$ is a Boolean automorphism. Boolean automorphisms map atoms into atoms. By Lemma 6.2, with $\tau=[i,j]$, $\mathsf{p}_{ij}^{\mathfrak{A}}$ is a bijection between atoms below $\mathsf{d}^{\mathfrak{A}}_{kl}$ and atoms below $\mathsf{d}^{\mathfrak{A}}_{\tau(k)\tau(l)}$. Clearly, the same is true for $\mathop{P}\nolimits_{ij}$ in place of $\mathsf{p}_{ij}$.
Take distinct atoms $a,b\in At$, $a,b\subseteq\mathsf{d}^{\mathfrak{A}}_{kl}$. By [12, 1.10.3], $\operatorname{C}_{k}^{U}(a)$ is an atom of the algebra of $\{k\}$-closed elements. In particular, $\operatorname{C}_{k}^{U}(a)$ and $\operatorname{C}_{k}^{U}(b)$ are either equal or disjoint. We claim first that $\operatorname{C}_{k}^{U}(a)=\operatorname{C}_{k}^{U}(b)$ is impossible. Clearly $\operatorname{C}_{k}^{U}(a)\neq\emptyset$, and $a=a\cap\mathsf{d}^{\mathfrak{A}}_{kl}$, $b=b\cap\mathsf{d}^{\mathfrak{A}}_{kl}$. If $\operatorname{C}_{k}^{U}(a)=\operatorname{C}_{k}^{U}(b)$, then
$\operatorname{C}_{k}^{U}(a\cap\mathsf{d}^{\mathfrak{A}}_{kl})\cap\operatorname{C}_{k}^{U}(b\cap\mathsf{d}^{\mathfrak{A}}_{kl})=\operatorname{C}_{k}^{U}(a)\neq\emptyset\,.$ (89)
On the other hand, $b\subseteq-a$, and thus (89) implies
$\operatorname{C}_{k}^{U}(a\cap\mathsf{d}^{\mathfrak{A}}_{kl})\cap\operatorname{C}_{k}^{U}(-a\cap\mathsf{d}^{\mathfrak{A}}_{kl})\neq\emptyset\,$, (90)
which contradicts cylindric axiom (C7). Therefore, we must have $\operatorname{C}_{k}^{U}(a)$ and $\operatorname{C}_{k}^{U}(b)$ disjoint.
By way of contradiction, suppose that there is $a\in At$, $a\subseteq\mathsf{d}^{\mathfrak{A}}_{kl}$ such that $\mathsf{p}_{ij}^{\mathfrak{A}}(a)\neq\mathop{P}\nolimits_{ij}^{U}(a)$. Then there must be an other atom $b\in At$, $b\subseteq\mathsf{d}^{\mathfrak{A}}_{kl}$ such that
$\mathsf{p}_{ij}^{\mathfrak{A}}(b)=\mathop{P}\nolimits_{ij}^{U}(a)\,.$ (91)
We already know that $\operatorname{C}_{k}^{U}(a)\neq\operatorname{C}_{k}^{U}(b)$ (in fact, they are disjoint), and automorphisms map disjoint elements into disjoint elements. In particular,
$\mathsf{p}_{ij}^{\mathfrak{A}}\big(\operatorname{C}_{k}^{U}(a)\big)\neq\mathsf{p}_{ij}^{\mathfrak{A}}\big(\operatorname{C}_{k}^{U}(b)\big)\,.$ (92)
Using that $\mathsf{p}_{ij}^{\mathfrak{A}}$ and $\mathop{P}\nolimits_{ij}^{U}$ agree on the $\operatorname{C}_{k}^{U}$-closed elements (Lemma 6.3) we get
[…]
But this contradicts (92). $\blacksquare$
Recall (see the paragraph before Lemma 4.4) that to every permutation $\sigma:\alpha\to\alpha$ there is a corresponding derived polyadic operation $\mathsf{p}_{\sigma}$ such that $\mathsf{p}_{[i,j]}=\mathsf{p}_{ij}$ and $\mathsf{p}_{\sigma\circ\tau}=\mathsf{p}_{\sigma}\circ\mathsf{p}_{\tau}$. The set-theoretic permutation operator is defined as
$\mathop{P}\nolimits_{\sigma}(x)=\big\{s\circ\sigma^{-1}:\;s\in x\big\}.$
It is easy to see that Lemmas 6.2, 6.3 and 6.4 extend for any permutation $\sigma$ in the following manner:
$\mathsf{p}_{\sigma}^{\mathfrak{A}}(\mathsf{d}_{kl}^{\mathfrak{A}})$ $=\mathsf{d}_{\sigma(k)\sigma(l)}^{\mathfrak{A}},$ (94)
$\mathsf{p}_{\sigma}^{\mathfrak{A}}(x)$ $=\mathop{P}\nolimits_{\sigma}^{U}(x)\quad\text{ for every }x\in B\,$, (95)
$\mathsf{p}_{\sigma}^{\mathfrak{A}}(a)$ $=\mathop{P}\nolimits_{\sigma}^{U}(a)\quad\text{ for every }a\in At,\ a\subseteq\mathsf{d}_{kl}^{\mathfrak{A}},\ k\neq l<\alpha\,.$ (96)
Let $PT_{\alpha}\subset T_{\alpha}$ denote the set of permutations of $\alpha$.
Call an atom $a\in A$ repetition-free if $a\not\subseteq\mathsf{d}_{kl}^{\mathfrak{A}}$ for every $k\neq l<\alpha$. Write
$RfAt=\big\{a\in At:\;(\forall k\neq l<\alpha)\ a\not\subseteq\mathsf{d}_{kl}^{\mathfrak{A}}\ \big\}\,$ (97)
and let $S$ denote the set of all sequences belonging to repetition-free atoms
$S$ $=\bigcup\big\{a:\;a\in RfAt\big\}\,.$ (98)
A sequence $s\in{}^{\alpha}U$ is called repetition-free if $s\notin\operatorname{D}_{kl}^{U}$ for every $k\neq l<\alpha$. Elements of $S$ are repetition-free sequences by (84), but $S$ is not necessarily the set of all such sequences, because atoms below some $\mathsf{d}_{kl}^{\mathfrak{A}}$ might also contain repetition-free sequences. We also have that $S$ is closed under permutations, i.e., for $s\in S$ and $\sigma\in PT_{\alpha}$, we have $s\circ\sigma\in S$. This is so, because $s\circ\sigma\notin S$ means that $s\circ\sigma\in a\subseteq\mathsf{d}_{kl}^{\mathfrak{A}}$ for some $k,l\in\alpha$. Then $s\in\mathop{P}\nolimits_{\sigma}(a)=\mathsf{p}_{\sigma}(a)\subseteq\mathsf{d}_{\sigma(k)\sigma(l)}^{\mathfrak{A}}$ by (96) and (94), so $s\notin S$ by the definition of $S$. For each $s\in S$ define its block as
$PT(s)=\{s\circ\sigma:\;\sigma\in PT_{\alpha}\}$ (99)
and let $R$ contain a representant element from each block, that is $|R\cap PT(s)|=1$ for each $s\in S$. Then
$S=\{s\circ\sigma:\;s\in R\mbox{ and }\sigma\in PT_{\alpha}\},$ (100)
and for each $z\in S$ there are unique $s\in R$ and $\sigma\in PT_{\alpha}$ such that $z=s\circ\sigma$.
Our task is to represent $\mathsf{p}_{\sigma}$ on the atoms by $\mathop{P}\nolimits_{\sigma}$. By (96), we have to change the representation only on the repetition-free atoms. How can it happen that $\mathsf{p}_{\sigma}(a)\neq\mathop{P}\nolimits_{\sigma}a$? For example, it may happen that $s\circ\sigma\in a\in RfAt$, then we would need $s\in\mathsf{p}_{\sigma}a$, but $s\in b\neq\mathsf{p}_{\sigma}a$. Instead of “pulling $s$ out of $b$ and putting it into $p_{\sigma}a$”, we make a “copy” $s_{\sigma}$ of $s$ and put $s_{\sigma}$ into $\mathsf{p}_{\sigma}(a)$. In more detail, we will make copies $s_{\sigma}$ of $s\in R$ for all $\sigma\in PT_{\alpha}$ and put $s_{\sigma}\circ\delta^{-1}$ to $\mathsf{p}_{\delta}\mathsf{p}_{\sigma}(a)$ whenever $s\circ\sigma\in a$. Then, as we will show, $\mathsf{p}_{\sigma}$ will be well-represented on $RfAt$ and the representations of the atoms will remain disjoint. How do we make copies $s_{\sigma}$ of $s$? First, we change $s$ into a set $\hat{s}$ of sequences as in the previous section by replacing $U$ with $U\times W$ for a suitable $W$, and then we will partition the set $\hat{s}$ to obtain the $s_{\sigma}$. We now begin to execute this plan.
[…]
$\hat{s}=\{s\times w:\;w\in{}^{\alpha}W\},$ (101)
and for $X\in\operatorname{Sb}({}^{\alpha}U)$ let
$F(X)=\bigcup\big\{\hat{s}:\;s\in X\big\}\in\operatorname{Sb}({}^{\alpha}V).$ (102)
It is straightforward to check that $F$ is an embedding
$\langle\operatorname{Sb}({}^{\alpha}U),\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U},\mathop{P}\nolimits_{ij}^{U}\,\rangle_{i<\alpha}\quad\hookrightarrow\quad\langle\operatorname{Sb}({}^{\alpha}V),\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathop{S}\nolimits_{ij}^{V},\mathop{P}\nolimits_{ij}^{V}\,\rangle_{i<\alpha}.$ (103)
Consequently, there is a subset $A^{\prime}\subseteq\operatorname{Sb}({}^{\alpha}V)$ such that $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is isomorphic to an algebra
$\mathfrak{A}^{\prime}=\langle A^{\prime},\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathop{S}\nolimits_{ij}^{V},\mathsf{p}_{ij}^{\prime},\mathsf{d}_{ij}^{\prime}\rangle_{i,j<\alpha},$ (104)
where $\mathsf{p}_{ij}^{\prime}$ and $\mathsf{d}_{ij}^{\prime}$ are the isomorphic images of $\mathsf{p}_{ij}^{\mathfrak{A}}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, respectively. Let $At^{\prime}$ denote the set of atoms of $\mathfrak{A}^{\prime}$, that is, the embedded image of $At$. Similarly, $\mathfrak{B}^{\prime}$ is the embedded image of $\mathfrak{B}$. For every permutation $\sigma$, equations (94)-(96) ensure
$\mathsf{p}_{\sigma}^{\prime}(\mathsf{d}_{kl}^{\prime})$ $=\mathsf{d}_{\sigma(k)\sigma(l)}^{\prime}$ (105)
$\mathsf{p}_{\sigma}^{\prime}(x)$ $=\mathop{P}\nolimits_{\sigma}^{V}(x)\quad\text{ for every }x\in B^{\prime}\,$, (106)
$\mathsf{p}_{\sigma}^{\prime}(a)$ $=\mathop{P}\nolimits_{\sigma}^{V}(a)\quad\text{ for every }a\in At^{\prime},\ a\subseteq\mathsf{d}_{kl}^{\prime},\ k\neq l<\alpha\,.$ (107)
Next, for each $s\in R$ we split the embedded image $\hat{s}$ ($F(\{s\})$) into $\alpha!$ ($\alpha$ factorial) elements.
Lemma 6.5.
[…]
Proof . Recall that $W=PT_{\alpha}$. Let $\star$ be any operation on $W$ such that $\langle W,\star\rangle$ is a commutative group, e.g., we can take $\star$ to be a bijective image of addition modulo $|W|$ on $\{0,1,\dots,|W|-1\}$. For $\sigma\in W$ define
$s_{\sigma}=\{s\times w\in{}^{\alpha}V:w_{1}\star\dots\star w_{n}=\sigma\}.$ (108)
It is now routine to check that $\{s_{\sigma}:\sigma\in PT_{\alpha}\}$ satisfies the requirements. $\blacksquare$
Let us change the representation $F(a)$ of the repetition-free atoms $a$ to $rep(a)$ defined as
$\operatorname{rep}(a)=\bigcup\big\{\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta}(a)\mbox{ and }s\in R\big\}.$ (109)
Lemma 6.6.
For $a\in RfAt$ and $\tau\in PT_{\alpha}$ we have
$\operatorname{rep}\big(\mathsf{p}_{\tau}^{\mathfrak{A}}(a)\big)=\mathop{P}\nolimits_{\tau}^{V}(\operatorname{rep}(a))\,.$ (110)
Proof of Lemma 6.6.
$\operatorname{rep}(\mathsf{p}_{\tau}^{\mathfrak{A}}(a))$ $=\bigcup\big\{\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta}(\mathsf{p}_{\tau}^{\mathfrak{A}}a),\ s\in R\big\}$ (111)
$=\bigcup\big\{\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta\circ\tau}(a),\ s\in R\big\}$ (112)
$=\bigcup\big\{\mathop{P}\nolimits_{\tau}\mathop{P}\nolimits_{\tau^{-1}}\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta\circ\tau}(a),\ s\in R\big\}$ (113)
$=\mathop{P}\nolimits_{\tau}\bigcup\big\{\mathop{P}\nolimits_{(\delta\circ\tau)^{-1}}(s_{\sigma}):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta\circ\tau}(a),\ s\in R\big\}$ (114)
$=\mathop{P}\nolimits_{\tau}(\operatorname{rep}(a)).$ (115)
$\blacksquare$
Recall that $F$ is defined in (102).
Lemma 6.7.
For $a\in RfAt$ and $i<\alpha$ we have
$\operatorname{C}_{i}^{V}(\operatorname{rep}(a))=\operatorname{C}_{i}^{V}(F(a))\,.$ (116)
Proof of Lemma 6.7. ($\supseteq$) Suppose $z\in\operatorname{C}_{i}^{V}(F(a))$ for some $a\in RfAt$ and $i<\alpha$. There are a repetition-free sequence $s\in R$ and a permutation $\sigma$ such that $s\circ\sigma\in a$, and $z\in\operatorname{C}_{i}^{V}(F(\{s\circ\sigma\}))$. With $\delta=\sigma$ in the definition (109) of $\operatorname{rep}(a)$, and using $s\circ\sigma\in a$ we get that $\mathop{P}\nolimits_{\sigma^{-1}}(s_{\sigma})\subseteq\operatorname{rep}(a)$. Observe that
$F(\{s\circ\sigma\})=\mathop{P}\nolimits_{\sigma^{-1}}(F(\{s\}))$ (117)
because $F$ is an embedding with respect to the set-theoretic polyadic operations. By construction, $s_{\sigma}\subseteq F(\{s\})$, and $\operatorname{C}_{j}^{V}(s_{\sigma})=\operatorname{C}_{j}^{V}(F(\{s\}))$ for all $j<\alpha$. Therefore,
[…]
$=$ $\mathop{P}\nolimits_{\sigma^{-1}}\operatorname{C}_{\sigma^{-1}(i)}^{V}(s_{\sigma})$ (120)
[…]
$\subseteq$ $\operatorname{C}_{i}^{V}(\operatorname{rep}(a)).$ (122)
($\subseteq$) In the other direction we need that $\operatorname{C}_{i}^{V}(\operatorname{rep}(a))\subseteq\operatorname{C}_{i}^{V}(F(a))$. By definition, and complete additivity of $\operatorname{C}_{i}^{V}$,
$\operatorname{C}_{i}^{V}(\operatorname{rep}(a))=\bigcup\big\{\operatorname{C}_{i}^{V}(\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma})):\;s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}\circ\delta}(a),\ s\in R\big\}.$ (123)
Suppose $z\in\operatorname{C}_{i}^{V}(\operatorname{rep}(a))$. Then there are $s\in R$, $\sigma,\delta\in PT_{\alpha}$ such that $s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}}\mathsf{p}^{\mathfrak{A}}_{\delta}(a)$, and $z\in\operatorname{C}_{i}^{V}(\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}))$. Now,
$\operatorname{C}_{i}^{V}(\mathop{P}\nolimits_{\delta^{-1}}(s_{\sigma}))=\operatorname{C}_{i}^{V}(\mathop{P}\nolimits_{\delta^{-1}}(F(\{s\})))$ (124)
(cf. steps (118)-(121)). Thus, $z\in\operatorname{C}_{i}^{V}(\mathop{P}\nolimits_{\delta^{-1}}(F(\{s\})))$. As $F$ is a homomorphism with respect to the “real” permutations,
$\mathop{P}\nolimits_{\delta^{-1}}(F(\{s\}))=F(\{s\circ\delta\}),$ (125)
and then $z\in\operatorname{C}_{i}^{V}(F(\{s\circ\delta\}))$.
To complete the proof, it is enough to show that $s\circ\sigma\in\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}}\mathsf{p}^{\mathfrak{A}}_{\delta}(a)$ implies $\operatorname{C}_{i}^{V}(F(\{s\circ\delta\}))\subseteq\operatorname{C}_{i}^{V}(F(a))$. Let $\eta=\delta^{-1}\circ\sigma$, and $b=\mathsf{p}^{\mathfrak{A}}_{\sigma^{-1}}\mathsf{p}^{\mathfrak{A}}_{\delta}(a)$. Now,
$s\circ\sigma\in b$ $\Longrightarrow$ $s\in\mathop{P}\nolimits_{\sigma}(b)$ (126)
$\Longrightarrow$ $s\circ\delta\in\mathop{P}\nolimits_{\delta^{-1}}\mathop{P}\nolimits_{\sigma}(b)=\mathop{P}\nolimits_{\eta}(b).$ (127)
Then
[…]
$\blacksquare$
We are ready to define our final representation
$G:\langle A,\cup,\smallsetminus,\operatorname{C}_{i}^{U},\mathop{S}\nolimits_{ij}^{U},\mathsf{p}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}\to\langle\operatorname{Sb}({}^{\alpha}V),\cup,\smallsetminus,\operatorname{C}_{i}^{V},\mathop{S}\nolimits_{ij}^{V},\mathop{P}\nolimits_{ij}^{V}\rangle_{i,j<\alpha}.$
For atoms $a\in At$ we define
$G(a)=\begin{cases}F(a)&\text{ if }a\subseteq\mathsf{d}_{kl}\text{ for some }k\neq l<\alpha,\\ \operatorname{rep}(a)&\text{ if }a\in RfAt\,.\end{cases}$ (130)
$G$ maps atoms of $\mathfrak{A}$ into elements of $\operatorname{Sb}({}^{\alpha}V)$. As $\mathfrak{A}$ is complete and atomic, $G$ extends to a Boolean embedding
$\langle A,\cup,\smallsetminus\rangle\quad\hookrightarrow\quad\langle\operatorname{Sb}({}^{\alpha}V),\cup,\smallsetminus\rangle\,$ (131)
by
$G(x)=\bigcup\{G(a):a\leq x,\ a\in At\}.$ (132)
Using Lemma 6.6 and (96), both for repetition-free and non repetition-free atoms $a$, $G(\mathsf{p}_{\sigma}^{\mathfrak{A}}a)=\mathop{P}\nolimits_{\sigma}(G(a))$, and thus $G$ is a homomorphism with respect to $\mathsf{p}$ as well.
$G$ is a homomorphism for the cylindrifications: we need that for any atom $a$, $G(\operatorname{C}_{i}^{U}(a))=\operatorname{C}_{i}^{V}(G(a))$. By Lemma 6.7, $\operatorname{C}_{i}^{V}(G(a))$ is always $\operatorname{C}_{i}^{V}(F(a))$. Now, if $z\in b$ for an atom $b\subseteq\operatorname{C}_{i}^{U}(a)$, then we have two cases. Either $b$ is not repetition-free, in which case $F(\{z\})\subseteq G(b)$; or $b$ is repetition-free. In this latter case there is $s\in R$ and $\tau\in PT_{\alpha}$ such that $z\circ\tau=s$, and then $z\circ\tau\in s_{\sigma}$ for some $\sigma$. This means that $F(\{z\})\subseteq\operatorname{rep}(c)$ for the atom $c$ for which $s\circ\sigma\in\mathsf{p}_{\sigma^{-1}}^{\mathfrak{A}}\mathsf{p}^{\mathfrak{A}}_{\tau^{-1}}(c)$. This $c$ must be below $\operatorname{C}_{i}^{U}(a)$. Putting the cases together, we end up at
$G(\operatorname{C}_{i}^{U}(a))=\bigcup\big\{G(b):\;b\subseteq\operatorname{C}_{i}^{U}(a),b\in At\big\}\supset\bigcup\big\{F(z):z\in\operatorname{C}_{i}^{U}(a)\big\}$ (133)
[…]
Finally, $G(\mathsf{d}_{ij}^{\mathfrak{A}})=F(\mathsf{d}_{ij}^{\mathfrak{A}})=\mathsf{d}_{ij}^{\prime}$. As $\mathop{S}\nolimits_{ij}^{V}$ can be expressed using $\operatorname{C}_{i}^{V}$ and $\mathsf{d}_{ij}^{\prime}$ (cf. (83)), it follows that $G(\mathop{S}\nolimits_{ij}^{U}(a))=\mathop{S}\nolimits_{ij}^{V}(G(a))$, and the proof is complete. $\blacksquare$
As a corollary to Theorem 6.1, we can fill in the status of the open line from $cs$ to cspcsp in [1, Figure 1]: it has to be a dashed line indicating finite axiomatizability.
Theorem 6.8.
$\mathsf{RA}_{\alpha}^{csp}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{cs}$ (by the equations (F0)–(F9) axiomatizing $\mathsf{FPA}_{\alpha}$).
Proof . This follows from (80) and $\mathsf{RA}_{\alpha}^{csp}\subseteq\mathsf{FPA}_{\alpha}$. $\blacksquare$
We now prove Theorem 1.3. The proof is completely analogous to that of Theorem 1.2. By $\mathsf{RFPA}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$ and $\operatorname{SRd}_{cs}\mathsf{RFPA}_{\alpha}=\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$, Theorem 1.3 states that $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{csp}$ iff ($\operatorname{Rd}_{cs}\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ and equations (10)–(13) hold in $\mathfrak{A}$). By Theorem 6.8 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{csp}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ and equations (F0)–(F9) hold in $\mathfrak{A}$). Equations (10)–(13) in Theorem 1.3 are exactly the equations in (F0)–(F9) in which the permutation operations occur. Theorem 1.3 has been proved.
We prove Theorem 1.1. By Theorem 1.2 and Theorem 1.3 we get that $\mathsf{RFPA}_{\alpha}$ is axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations (F0)–(F9) defining $\mathsf{FPA}_{\alpha}$. [22, Theorem 1(i)] states that $\mathsf{FPA}_{\alpha}$ is term-definitionally equivalent to $\mathsf{PA}_{\alpha}$. The same term-definitions take $\mathsf{RFPA}_{\alpha}$ to $\mathsf{RPA}_{\alpha}$. Hence, $\mathsf{RPA}_{\alpha}$ is finitely axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations defining $\mathsf{PA}_{\alpha}$. Examining the defining equations of $\mathsf{PA}$, in e.g., [12, Def.5.4.1], one sees that among these equations (1)–(4) are the ones in which the substitution operations $\mathsf{s}_{\sigma}$ occur and the rest follow easily from (F0)–(F9). Theorem 1.1 has been proved.
We conclude by recalling a question from [1], which remains open.
Problem 6.9.
Are permutations finitely axiomatizable over the Boolean operations together with cylindrifications? I.e., does the class $\mathsf{RA}^{cp}_{\alpha}$ admit a finite axiomatization over $\mathsf{RA}^{c}_{\alpha}$?

### v2

@@
 $\mathsf{s}_{ij}(x+y)=\mathsf{s}_{ij}(x)+\mathsf{s}_{ij}(y),\quad\mathsf{s}_{ij}(-x)=-\mathsf{s}_{ij}(x),$ (6)
-$\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x,\quad\mathsf{s}_{ii}(x)=x,$ (7)
+$\mathsf{s}_{ij}\mathsf{c}_{i}x=\mathsf{c}_{i}x,\quad\mathsf{s}_{ij}x=\mathsf{c}_{i}\mathsf{s}_{ij}x,\quad\mathsf{s}_{ii}(x)=x,$ (7)
 $\mathsf{s}_{ij}\mathsf{c}_{k}x=\mathsf{c}_{k}\mathsf{s}_{ij}x,$ (8)
-$\mathsf{s}_{ij}x=\mathsf{c}_{i}\mathsf{s}_{ij}x.$ (9)
+$\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x.$ (9)
 Theorem 1.3 ($\mathsf{p}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i},\mathsf{s}_{ij}$).
@@
 $\mathsf{p}_{ij}(x+y)=\mathsf{p}_{ij}(x)+\mathsf{p}_{ij}(y),\quad\mathsf{p}_{ij}(-x)=-\mathsf{p}_{ij}(x),$ (10)
-$\mathsf{p}_{ij}\mathsf{p}_{ij}x=x,\quad\mathsf{p}_{ii}x=x,\quad\mathsf{p}_{ij}x=\mathsf{p}_{ji}x,$ (11)
-$\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x,$ (12)
-$\mathsf{p}_{ij}\mathsf{s}_{ij}x=\mathsf{s}_{ji}\mathsf{p}_{ij}x.$ (13)
+$\mathsf{p}_{ij}\mathsf{p}_{ij}x=x,\quad\mathsf{p}_{ij}x=\mathsf{p}_{ji}x,\quad\mathsf{p}_{ii}x=x,$ (11)
+$\mathsf{p}_{ij}\mathsf{s}_{ij}x=\mathsf{s}_{ji}\mathsf{p}_{ij}x,$ (12)
+$\mathsf{p}_{ij}\mathsf{p}_{ik}x=\mathsf{p}_{jk}\mathsf{p}_{ij}x.$ (13)
 We do not know whether $\mathsf{p}_{ij}$ is finitely axiomatizable over $\mathsf{c}_{i}$, but we know that $\mathsf{s}_{ij}$ is finitely axiomatisable over $\mathsf{c}_{i},\mathsf{p}_{ij}$, see [3].
@@
 The equation $\mathsf{s}_{ij}(x)=\mathsf{s}_{ij}(-x)$ is part of what it means that $\mathsf{s}_{ij}$ is a Boolean endomorphism, and it is visible that the negation symbol is used here.
-This equation, as noted in [22, p.561], can be replaced by the ones expressing that $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$.
+This equation, as noted in [23, p.561], can be replaced by the ones expressing that $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$.
 This is because if $\mathsf{s}_{ij}$ preserves ++, $\cdot$, 0 and $1$, then
@@
 and these imply $\mathsf{s}_{ij}(-x)=-\mathsf{s}_{ij}(x)$. The same comment applies to the case of $\mathsf{p}_{ij}$. Hence, $\mathsf{FPA}_{\alpha}$ is a canonical variety.
-In Section 6, we will use the following consequences of the $\mathsf{FPA}_{\alpha}$-axioms.
+In Section 6, we will use consequences (P1)(P5) of the $\mathsf{FPA}_{\alpha}$-axioms.
+In Definition 4.5 below and in section 5 we will use consequence (FS) of the $\mathsf{FPA}_{\alpha}$-axioms that concerns solely the substitution operations $\mathsf{s}_{ij},i,j<\alpha$.
+It is proved in [23, Remark 7] that (FS) does not follow from the axioms of $\mathsf{FPA}_{\alpha}$ that do not contain any reference to $\mathsf{p}_{ij},i,j<\alpha$.
 Lemma 4.3.
@@
     $\mathsf{p}_{ij}\mathsf{c}_{m}x=\mathsf{c}_{m}\mathsf{p}_{ij}x$ if $i,j,m<\alpha$ are distinct.
+  6. (FS)
+    $\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x$ if $i,j,k<\alpha$ are distinct.
 Proof . Here are the derivations.
@@
 However, the proof of (the typo-free version of) Lemma 1.6 in [23] is correct, and throughout that paper the lemma is used with the correct statement.
+(FS) is proved as Lemma 1.5(iii) in [23].
 $\blacksquare$
@@
 This way to every permutation $\sigma$ we assign a corresponding polyadic term $\mathsf{p}_{\sigma}$.
-[22, Claim 1.2] shows that different decompositions of the same permutation into products of transpositions yield identical terms (in $\mathsf{FPA}_{\alpha}$).
-For a precise definition we refer to [22, p.554].
+[23, Claim 1.2] shows that different decompositions of the same permutation into products of transpositions yield identical terms (in $\mathsf{FPA}_{\alpha}$).
+For a precise definition we refer to [23, p.554].
 This leads to the important lemma below that we make use of many times in Section 6.
@@
 Next we turn to algebras without the $\mathsf{p}_{ij}$ operations.
-The class $\mathsf{SCA}_{\alpha}$ of algebras is axiomatized by the finitary polyadic algebra axioms of $\mathsf{FPA}_{\alpha}$ without the $\mathsf{p}_{ij}$’s, that is, when the $\mathsf{FPA}_{\alpha}$ axioms speaking about $\mathsf{p}_{ij}$ are deleted.
+The class $\mathsf{SCA}_{\alpha}$ of algebras is axiomatized by the finitary polyadic algebra axioms of $\mathsf{FPA}_{\alpha}$ without the $\mathsf{p}_{ij}$’s, that is, when the $\mathsf{FPA}_{\alpha}$ axioms speaking about $\mathsf{p}_{ij}$ are deleted; with (FS) added as axiom.
 We will see that these algebras are obtained from cylindric algebras by replacing the diagonal constants $\mathsf{d}_{ij}$ with the derived operations $\mathsf{s}_{ij}$.
 The elements of $\mathsf{SCA}_{\alpha}$ will be called substitution-cylindric algebras.
-Definition 4.5 (Cf. Sax on p.559 in [22]).
+This class was introduced by Pinter [20] where it is denoted by $\mathsf{QA}_{\alpha}$, and its members are called quantifier-algebras.
+Definition 4.5 (Cf. Definition 2.1 in [20]).
 The class $\mathsf{SCA}_{\alpha}$ consists of the algebras $\mathfrak{A}=\langle A,+,-,\mathsf{c}_{i},\mathsf{s}_{ij}\rangle_{i,j<\alpha}$ in which the equations below are valid for every $i,j,k<\alpha$.
@@
     $\mathsf{s}_{ij}$ are Boolean endomorphisms
+  8. (FS)
+    $\mathsf{s}_{ij}\mathsf{s}_{kj}x=\mathsf{s}_{ij}\mathsf{s}_{ki}x$ if $i,j,k<\alpha$ are distinct.
 We deliberately use the same labels as in Definition 4.1, even if (F0) and (F6) here are not exactly the same as (F0) and (F6) in Definition 4.1 as here the $\mathsf{p}_{ij}$’s are missing.
@@
 The proof of (i) relies on results in Pinter [19] and Sain-Thompson [23].
-First we prove that (F0)-(F6) imply substitution axioms (S1)-(S6) in [19].
-Indeed, Lemma 1.5(iii) and Proposition 6 in [22] prove that the equations
-  1. (A1)
-    $\mathsf{s}_{ij}\mathsf{s}_{ki}x=\mathsf{s}_{ij}\mathsf{s}_{kj}x$
+First we prove that (F0)-(F6),(FS) imply substitution axioms (S1)-(S6) in [19].
+Indeed, Proposition 6 in [23] proves that the equations
   1. (A1)
@@
     $\mathsf{s}_{ij}\mathsf{s}_{kl}x=\mathsf{s}_{kl}\mathsf{s}_{ij}x$ if $i\notin\{k,l\}$, $k\neq j$
-hold in $\mathsf{SCA}_{\alpha}$ and (S1)-(S6) are (F6), (F0), (A1), (A2), and (A3). Also, [22] proves
+hold in $\mathsf{SCA}_{\alpha}$ and (S1)-(S6) are (F6), (F0), (FS),(A1), and (A2). Also, [23] proves
 $\mathsf{c}_{i}x$ is the least element of $\{y\in\operatorname{ran}\mathsf{s}_{ij}\}$
@@
 To prove (ii), notice first that $\mathsf{SCA}_{\alpha}\subseteq\mathbf{S}\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$ follows from (i) together with $\mathsf{SCA}_{\alpha}$ being a canonical variety.
-For the other direction, we have to prove that (F0)-(F6) hold in $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$.
-Now, (F0) is (C0)+[12, 1.5.1], (F1) is (C2), and (F2)-(F6) are [12, 1.2.6(ii), 1.5.8(i), 1.5.9(ii), 1.5.8(ii), 1.5.3], respectively.
+For the other direction, we have to prove that (F0)-(F6),(FS) hold in $\operatorname{Rd}_{sc}\mathsf{CA}_{\alpha}$.
+Now, (F0) is (C0)+[12, 1.5.1], (F1) is (C2), and (F2)-(F6),(FS) are [12, 1.2.6(ii), 1.5.8(i), 1.5.9(ii), 1.5.8(ii), 1.5.3, 1.5.10(ii)], respectively.
 $\blacksquare$
@@
 Theorem 5.10.
-$\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$ (by the equations (F0)-(F6) defining $\mathsf{SCA}_{\alpha}$).
+$\mathsf{RA}_{\alpha}^{cs}$ is finitely axiomatizable over $\mathsf{RA}_{\alpha}^{c}$ (by the equations (F0)-(F6),(FS) defining $\mathsf{SCA}_{\alpha}$).
 Proof . This is an immediate corollary of Lemma 5.2 and $\mathsf{RA}_{\alpha}^{cs}\subseteq\mathsf{SCA}_{\alpha}$. $\blacksquare$
@@
 So Theorem 1.2 states that $\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{cs}$ iff ($\operatorname{Rd}_{c}\mathfrak{A}\in\mathbf{SP}\mathsf{RA}_{\alpha}^{c}$ and equations (6)–(9) hold in $\mathfrak{A}$).
-By Theorem 5.10 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{c}$ and equations (F0)-(F6) hold in $\mathfrak{A}$).
-Equations (6)–(9) in Theorem 1.2 are the equations in (F0)-(F6) in which the substitutions occur [^11] The indices $i,j,k$ in Theorem 1.2 are required to be distinct while in (F0)–(F6) they are not required to be distinct.
+By Theorem 5.10 we get that $\mathfrak{A}\in\mathsf{RA}_{\alpha}^{cs}$ iff ($\mathfrak{A}\in\mathsf{RA}_{\alpha}^{c}$ and equations (F0)-(F6),(FS) hold in $\mathfrak{A}$).
+Equations (6)–(9) in Theorem 1.2 are the equations in (F0)-(F6),(FS) in which the substitutions occur [^11] The indices $i,j,k$ in Theorem 1.2 are required to be distinct while in (F0)–(F6),(FS) they are not required to be distinct.
 It is easy to check that this does not make any difference.; the ones in which the substitutions do not occur hold in $\mathsf{RA}_{\alpha}^{c}$, so they are superfluous.
@@
 For any $i,j<\alpha$, from (83), [12, 1.5.4(i)] and cylindric axiom (C2) we get $\mathop{S}\nolimits_{ij}^{U}(\mathsf{d}_{ij}^{\mathfrak{A}})=1$ and $x\cap\mathsf{d}_{ij}^{\mathfrak{A}}\subseteq\mathop{S}\nolimits_{ij}^{U}(x)$.
-Sain–Thompson [22] define the class $\mathsf{FPEA}_{\alpha}$ as $\mathsf{FPA}_{\alpha}$ enriched with diagonal elements $\mathsf{d}_{ij}$ satisfying $\mathsf{s}_{ij}\mathsf{d}_{ij}=1$ and $x\cdot\mathsf{d}_{ij}\leq\mathsf{s}_{ij}x$ (see [22, p.544, Definition 2]).
+Sain–Thompson [23] define the class $\mathsf{FPEA}_{\alpha}$ as $\mathsf{FPA}_{\alpha}$ enriched with diagonal elements $\mathsf{d}_{ij}$ satisfying $\mathsf{s}_{ij}\mathsf{d}_{ij}=1$ and $x\cdot\mathsf{d}_{ij}\leq\mathsf{s}_{ij}x$ (see [23, p.544, Definition 2]).
 We just verified that these two properties hold for $\mathop{S}\nolimits_{ij}^{U}$ and $\mathsf{d}_{ij}^{\mathfrak{A}}$, therefore the expanded structure $\langle\mathfrak{A},\mathsf{d}_{ij}^{\mathfrak{A}}\rangle_{i,j<\alpha}$ is an $\mathsf{FPEA}_{\alpha}$.
-[22] proves that $\mathsf{FPEA}_{\alpha}$ are term equivalent to quasi-polyadic equality algebras $\mathsf{QPEA}_{\alpha}$, and in particular, [22, p.557-558, (E3)] proves that from $\mathsf{FPEA}_{\alpha}$ the identity
+[23] proves that $\mathsf{FPEA}_{\alpha}$ are term equivalent to quasi-polyadic equality algebras $\mathsf{QPEA}_{\alpha}$, and in particular, [23, p.557-558, (E3)] proves that from $\mathsf{FPEA}_{\alpha}$ the identity
 $\mathsf{p}_{ij}(\mathsf{d}_{kl})=\mathsf{d}_{\tau(k)\tau(l)}$ (85)
@@
 We prove Theorem 1.1.
-By Theorem 1.2 and Theorem 1.3 we get that $\mathsf{RFPA}_{\alpha}$ is axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations (F0)–(F9) defining $\mathsf{FPA}_{\alpha}$.
-[22, Theorem 1(i)] states that $\mathsf{FPA}_{\alpha}$ is term-definitionally equivalent to $\mathsf{PA}_{\alpha}$.
+By Theorems 1.21.3 and the last item (FS) in Lemma 4.3 we get that $\mathsf{RFPA}_{\alpha}$ is axiomatized over $\mathsf{Rdf}_{\alpha}$ by the equations (F0)–(F9) defining $\mathsf{FPA}_{\alpha}$.
+[23, Theorem 1(i)] states that $\mathsf{FPA}_{\alpha}$ is term-definitionally equivalent to $\mathsf{PA}_{\alpha}$.
 The same term-definitions take $\mathsf{RFPA}_{\alpha}$ to $\mathsf{RPA}_{\alpha}$.